### v1

# A Gentle Wakeup Call: Symmetry Breaking with Less Collision Cost

Abstract
The wakeup problem addresses the fundamental challenge of symmetry breaking. There are $n$ devices sharing a time-slotted multiple access channel. In any fixed slot, if a single device sends a packet, it succeeds; however, if two or more devices send, then there is a collision and none of the corresponding packets succeed. For the static version of wakeup, all packets are initially active (i.e., can send and listen on the channel); for the dynamic version, the packets become active at arbitrary times. In both versions, the goal is to successfully send a single packet.
Prior results on wakeup have largely focused on the number of slots until the first success; that is, the latency. However, in many modern systems, collisions introduce significant delay, an aspect that current wakeup algorithms do not address. For instance, while existing results for static wakeup have polylogarithmic-in-$n$ latency, they can incur additional latency that is linear in the cost of a collision $\mathcal{C}$. Thus, the total latency is large and dominated by the contributions from collisions.
Here, we design and analyze a randomized wakeup algorithm, Aim-High. For sufficiently large $\mathcal{C}$ and with bounded error, Aim-High has latency and expected collision cost that is nearly $O(\sqrt{\mathcal{C}})$ for both the static and dynamic versions. Otherwise, the latency and expected collision cost are $O({\texttt{poly}({\log n})})$ for the static setting, and $O(n\,{\texttt{poly}({\log n})})$ for the dynamic setting. We also establish lower bounds that complement these results.
Keywords:
Wakeup, symmetry breaking, multiple access channel, collision cost
1 Introduction
In the wakeup problem, devices are classified as either active (awake) or inactive (asleep). Each active device has a packet to send on a shared channel, while inactive participants are limited to listening. The first successful transmission acts as a “wakeup call” where all sleeping participants are woken up, and the problem is solved [2].
Traditionally, prior work on this problem has focused on the time until the first success, which is typically referred to as the latency. Minimizing this metric is challenging, since in any unit of time, the shared channel can only accommodate a single sender; if two or more devices send simultaneously, the result is a collision where none of the corresponding packets are successfully sent. Moreover, sending on the channel is performed in a distributed fashion; that is, a priori there is no central scheduler or coordinator.
Under the standard cost model for wakeup, each collision wastes a single slot. However, in many settings, a collision can cost us much more than a single slot. For instance, in mobile systems, a failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range [26]. For WiFi networks, a collision is known to consume time proportional to the packet size [1]. Therefore, collisions may significantly impact the performance of modern systems.
In this work, we address both latency and collision cost in solving the wakeup problem. As we will discuss later in Section 3.1, while existing algorithms and natural approaches yield strong latency guarantees, they often result in high collision costs. Therefore, optimizing for these two metrics requires a new approach.
1.1 Model and Notation
[…]
Communication. Time is divided into disjoint slots, each of which can accommodate a packet transmission. Communication occurs on a multiple access channel, which is defined as follows. In any slot, a packet can either attempt to transmit or listen to the channel. If no packet is transmitted in a given slot, it is empty. If only one packet is transmitted, the packet succeeds; we often refer to this as a success. However, if multiple packets transmit simultaneously, they all fail—this is a collision—and they may try again later. There is no ability to distinguish between an empty slot and a collision; that is, there is no collision detection. There are no identifiers for the packets. Finally, packets must be sent in a distributed manner; that is, there is no scheduler or central authority.
Packet Activation. We consider two settings. For the static setting, in the first slot, all $n$ packets are active on the multiple access channel and begin executing an algorithm. By implication, all packets agree on the index of the current slot; this feature is typically referred to as a global clock [2].
In the dynamic setting, all $n$ packets become active over time. To capture the worst case, we consider these activation times to be scheduled by an adversary, and we often refer to packets being activated or injected by the adversary. Packets may not agree on the index of the current slot, and thus there is no global clock.
Metrics. We care about latency, which refers to the time between the first packet injection and the first success. Additionally, each collision incurs a known cost of $\mathcal{C}$ $\geq 4$. [^1] The value $4$ is chosen to simplify the analysis. We also note that only $\mathcal{C}$ $\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [16, 13, 2]) for the standard model will be asymptotically optimal. To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown. Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
Given any wakeup algorithm, the pertinent costs are: (i) the collision cost, which is the number of collisions multiplied by $\mathcal{C}$, and (ii) the latency. Although both metrics represent time-based costs and we aim to minimize their maximum, we report them separately for clarity and to facilitate our lower-bound results.
[…]
For a random variable $X$, we say that w.h.p. in $x$ the expected value $E[X]\leq v$ if, when tallying the events in the expected value calculation, the total probability of those events where $X>v$ is $O(1/{\texttt{poly}({x})})$, where the constant degree of the polynomial is can be made arbitrarily large.
Finally, our results are parameterized by $\epsilon$, which is a small positive constant $\ll 1$ set by our algorithm prior to execution. For terms of the form $\log^{k+(k^{\prime}/\epsilon)}n$, where $k,k^{\prime}$ are positive constants at least $1$, we report this as $\log^{\Theta(1/\epsilon)}n$ for ease of presentation.
1.2 Our Results
Our algorithm, Aim-High (AH), applies to the static and dynamic wakeup problems. [^2] We explain the name in Section 3.1 We first address the static setting, which is closely tied to the classic problem of leader election (see Section 2). Here, we define $\eta$ $\max\{n,\mathcal{C}\}$, which is used when bounding the error of our upper bounds derived in Section 4.1.
Theorem 1.1
W.h.p. in $\eta$, AH solves the static wakeup problem. The following associated costs hold w.h.p. in $\eta$:
  1. 1.
    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+2\epsilon}\right)$.
  2. 2.
    Else, the latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
By way of comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1). In contrast, Theorem 1.1 implies that if $\mathcal{C}$ is “large”, AH offers performance that is significantly sublinear in $\mathcal{C}$. Otherwise, the latency and the expected collision cost are each polylogarithmic-in-$n$, which echoes prior wakeup results without collision detection (see Section 2).
Many upper bounds in the static setting are achieved by “fair” algorithms. A fair algorithm is defined by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability used by every packet in the $i$-th slot. In Section 4.2, we give the following lower bound for this class of algorithms (to which AH belongs) having a w.h.p. bound on the expected latency.
Theorem 1.2
Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$. Any fair algorithm that w.h.p. in $\eta$ solves the static wakeup problem with $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has w.h.p. an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-k\epsilon}/\log^{\beta}n)$.
This theorem, along with noting that the bound in (2) of Theorem 1.1 is within a polylogarithmic-in-$n$ factor of the optimal latency and expected collision cost, yields the following result.
Corollary 1
When $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, AH is within $O(\mathcal{C}^{4\epsilon})$-factor of the optimal expected collision cost for fair algorithms, and within a polylogarithmic-in-$n$ factor of the optimal latency and expected collision cost otherwise.
In Section 4.3, we note that while AH upper bounds the latency and expected collision cost with w.h.p. in $\eta$, if this guarantee fails, then these quantities may be unbounded. We describe revisions to AH that yield the same bounds for expected latency and expected collision cost that hold with probability $1$.
Next, we analyze the performance of AH in the dynamic setting, and we give the following upper-bound results. Here, our w.h.p. guarantees are in terms of $\mathcal{C}$.
Theorem 1.3
W.h.p. in $\mathcal{C}$, AH solves the dynamic wakeup problem. The following associated costs hold w.h.p. in $\mathcal{C}$:
  1. 1.
    If  $\mathcal{C}=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$ for a small constant $\delta>0$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+\epsilon}\log n\right)$,
  2. 2.
    Else, the latency is $\tilde{O}(n(\log^{\Theta(1/\epsilon)}n))$ and the expected collision cost is $\tilde{O}(n(\log^{\Theta(1/\epsilon)}n))$.
Again, if $\mathcal{C}$ is large, then we achieve performance that is sublinear in $\mathcal{C}$. We give a lower bound for a class of algorithms we call “batch-fair" (to which AH belongs) that is analogous to fair algorithms, but natural in the dynamic setting.
Theorem 1.4
When $\mathcal{C}=\Omega(n^{2-\delta})$ for a small constant $\delta>0$, AH is within an $O(\mathcal{C}^{2\epsilon}\log^{2}n)$-factor of the optimal expected collision cost for batch-fair algorithms, and within an $O(\log^{\Theta(1/\epsilon)}n)$-factor of the optimal latency otherwise.
2 Related Work
The wakeup problem, has been extensively studied for decades. Here, we summarize closely-related results.
Static Wakeup. In the literature, static wakeup, also known as leader election, addresses the fundamental problem of symmetry breaking. Early work by Willard [25] gives an upper bound of $\lg\lg n+O(1)$, and a matching lower bound for fair algorithms with collision detection (CD).
In contrast, without CD, a common theme is that the latency increases to logarithmic or polylogarithmic in $n$. For starters, without collision detection and when $n$ is known, Kushilevitz and Mansour [21, 20] prove a lower bound of $\Omega(\log n)$ expected latency. Newport [23, 24], in work that unifies many prior lower-bound results, shows $\Omega(\log n)$ expected latency and $\Omega(\log^{2}n)$ latency w.h.p. for general randomized algorithms. Complementing these lower bounds, Bar-Yehuda et al. [3] achieve a matching upper bound of $O(\log n)$ expected latency and w.h.p. $O(\log^{2}n)$ latency.
Farach-Colton et al. [10] obtain $\Omega(\log n\log(1/\varepsilon))$ latency for algorithms without CD, succeeding with probability at least $1-\varepsilon$. A complementary by Jurdziński and Stachowiak [18] demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
Given access to $b>1$ multiple channels with CD, Fineman et al. [12] give an algorithm with $O((\log n/\log b)+(\log\log n)(\log\log\log n))$ latency whp for any number of active devices. In the SINR model [22], Fineman et al. [11] demonstrate $O(\log n+\log R)$ latency w.h.p., where $R$ is the ratio of longest to shortest link, implying $O(\log n)$ latency as typically $R=O({\texttt{poly}({n})})$; they also provide a general $\Omega(\log n)$ lower bound. Chang et al [7] focus on energy efficiency under different models of channel feedback. Gilbert et al. [14] analyze the impact of system-size predictions, deriving lower and upper bounds.
Dynamic Wakeup. The static and dynamic wakeup problems differ significantly. Specifically, for dynamic wakeup without CD, Jurdzinski and Stachowiak [19] show an $\Omega(\log^{2}n)$ lower bound, which contrasts with the $\Omega(\log n)$ lower bound [21] and matching upper bound [3] in the static setting.
Without CD, Jurdziński and Stachowiak [18] explore the implications of global and local clocks, known or unknown $n$, packet labels, both with and without CD. Most relevant to our work is an algorithm that succeeds with probabilty at least $1-\varepsilon$ and has $O(n/\log n)$ latency when there is no knowledge of $n$, no global clock, no CD, and no labels; the probability of failure for the corresponding algorithm is parameterized
Under the SINR model without a global clock, assuming devices know a polynomial upper bound on $n$, Jurdzinski and Stachowiak [19] give a Monte Carlo algorithm with $O(\log^{2}n/\log\log n)$ latency w.h.p., and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency, compared to a deterministic $O(\log^{2}n)$ latency [17]. Without CD and no global clock, Chlebus et al. [9] explore the impact of multiple $b>1$ channels, where they give a randomized algorithm with error $\varepsilon$ and latency $O(n^{1/b}\ln(1/\varepsilon))$.
Finally, we highlight prior work that addresses the cost of collisions in the context of the closely-related problem of contention resolution (see [8, 2]), where the goal is to have all $n$ packets successfully sent . Here, Biswas et al. [5] solve the static setting of contention resolution with latency and expected collision cost $\tilde{O}(n\sqrt{\mathcal{C}})$; however, they do not address the dynamic setting. Furthermore, Biswas et al. [5] assume CD, which their algorithm relies on to adjust the sending probability; this additional channel feedback is unavailable in our model. Another related result is by Anderton et al. [1] who show that collisions significantly degrade the performance of popular backoff algorithms for contention resolution.
3 Algorithm Specification
1
2
3// Initialization
4 $w_{0}\leftarrow 2^{\mathcal{C}^{\epsilon}}$
5 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
6 // Halving Phase
7 while $w_{\mbox{\tiny cur}}\geq 2$ do
8 for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ell$ do
9
10   Send packet with probability $1/w_{cur}$
11 if success then
12     Terminate
13 else
14    $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$
15
16// Doubling Phase
17 while True do
18 $w_{\mbox{\tiny cur}}\leftarrow 2w_{\mbox{\tiny cur}}$
19 for slot $i=1$ to $d\ell$ do
20
21   Send packet with probability $1/w_{cur}$
22 if success then
23     Terminate
24
Algorithm 1 Aim-High
The pseudocode for AH is given above, and each packet executes this algorithm upon being activated. Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$; recall that $\epsilon$ $\ll 1$ is a small positive constant that influences our bounds on cost.
For the static setting, $\ell$ $\ln(w_{\mbox{\tiny cur}})$, where $w_{\mbox{\tiny cur}}$ is the current window size. For the dynamic setting, $\ell=\ln(\mathcal{C})$.
Initially, the current window size $w_{\mbox{\tiny cur}}$ is set to $w_{0}$, and this is updated throughout the execution of two phases, as we describe below and illustrated in Figure 1.
[…]
After finishing a sample, if no success has occurred, then each packet sets $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$. This is repeated until either a success (and termination) occurs, or further halving would reduce the window size below $2$ (Line 1). At which point, the doubling phase begins.
Doubling phase. At the beginning of this phase, the window size doubles such that now $w_{\mbox{\tiny cur}}=4$ (Line 1). Then, each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of $d\ell$ slots (Line 1 and 1); we refer to these slots as a small sample. If a success occurs, the execution terminates; otherwise, this process continues.
3.1 Context and Overview
In this section, we start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
3.1.1 A Baseline via Backoff.
In the static setting, a natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$. At $i=\lg(n)$, the probability of a success in a slot becomes constant. This approach is a common ingredient in algorithm design; indeed, AH makes use of it. Another example of its use in solving wakeup is by Jurdziński and Stachowiak [18], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours, but without a notion of collision cost. However, in this prior result, each active packet starts with a constant sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
So, we can observe that while backoff may be an important part of any solution, a key obstacle is how to set the initial sending probability suitably low. We do not know $n$ a priori, and so it seems that no matter what initial sending probability we choose—even if it depends on the known value of $\mathcal{C}$—the adversary may choose a sufficiently large value of $n$ such that collisions are unavoidable. What can we do?
Figure 1: Illustration of the phases for AH. The top portion depicts the halving phase, where the initial window size is $w_{0}=2^{\mathcal{C}^{\epsilon}}$. The bottom portion depicts the doubling phase, starting from a window of size $4$, where the degree to which packets contend for channel access decreases as the window grows, leading to a likely success.
3.1.2 Design Choices and Overview.
We now discuss how the design of AH allows us to do better than prior results in a model with collision cost.
Setting an Initial Window Size. The initial sending probability needs to be set carefully. Too high and we incur many (potentially expensive) collisions as we have seen above.
Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as a good window. Informally, once packets reach this good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$. Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success. At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots. This observation dictates the size of our large samples (recall Section 3).
In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above informal analysis yields good expected collision cost (and, as we will show, good latency). However, since the adversary sets $\mathcal{C}$ with knowledge of $n$, this is by no means guaranteed; the adversary can always make collisions overwhelmingly likely. However, in this case, our consolation prize is that collisions should be cheap. To elaborate, we choose to "aim high" with the initial window size; that is, setting $w_{0}=2^{\mathcal{C}^{\epsilon}}$. This choice sets up our analysis of the following two cases for the static setting:
  1. 1.
    Go Left For Good Window. If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good window exists at or to the left of the initial window; that is, the good-window size is less than or equal to the initial-window size. Here, the probability of a collision is “not too high”. Thus, so long as we adjust the sending probability upwards in a conservative manner, we can reduce expected collision cost; this is the purpose of the halving phase. Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the time until the first success will be too large.
  2. 2.
    Go Right For Good Window Conversely, if $n\sqrt{\mathcal{C}}>w_{0}$, then the good window exists to the right of the initial window. Here, the collision probability may be “too high”. However, a key realization here is that the collision cost is cheap, since we will argue that $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$ implies that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, we can tolerate these collisions as we increase the window size in the doubling phase.
While the above cases address the particulars of our approach to the static setting, analogous intuition holds for the dynamic version (see Section 5.1 for discussion).
Samples. As discussed above, we "aim high" with the initial window and execute the the halving phase; here, we are hoping to succeed via samples of size $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$. However, if the adversary chooses to set $\mathcal{C}$ such that $n\sqrt{\mathcal{C}}>w_{0}$, good contention will not be discovered by the halving phase, and our ambitious setting of $w_{0}$ does not yield a success in the halving phase.
In this case, however, collisions are now provably cheap, and we no longer care about finding a good window; we simply want a quick success, even if we rack up collisions in the process. Thus, the doubling phase—which is classic backoff—uses smaller samples of size $d\log(w_{\mbox{\tiny cur}})$, and once a window of size $\Theta(n)$ is reached, a success is likely.
Finally, we note that in the static setting, the start and end points of each sample are the same for all packets. However, for the dynamic setting, this may not hold given that the adversary controls when packets are activated; that is, the samples of packets can be unsynchronized, and this poses an additional challenge.
Parameter $\epsilon$. Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis. Yet, it is worth pointing out how it interacts with our metrics of interest: latency and expected collision cost. The former is improved by making $\epsilon$ small, since this shortens the halving phase. So, if aiming high pays off, then a smaller $\epsilon$ can reduce latency. However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, degree of the polynomial increases with the reciprocal of $\epsilon$, which means our expected collision cost increases.
4 The Static Setting
In this section, we analyze the performance of AH for the static wakeup problem. Throughout our analysis, due to space constraints, some proofs have been moved to the appendix.
4.1 Upper Bound
Throughout, we assume that $n$ is sufficiently large. We make use of the following well-known facts.
Fact 1
The following inequalities hold.
  1. (a)
    For any $x$, $1-x\leq e^{-x}$,
  2. (b)
    For any $0\leq x\leq 1/2$, $1-x\geq e^{-2x}$.
We also make use of the following standard Chernoff bounds.
Theorem 4.1
(Goemans [15]) Let $X=\sum_{i}X_{i}$, where $X_{i}$ are Bernoulli indicator random variables with probability $p$. The following holds:
$Pr\left(X\geq(1+\delta)E[X]\right)$ $\leq\exp\left(-\frac{\delta^{2}E[X]}{2+\delta}\right)\mbox{~~~for any $\delta>0$}$
and
[…]
4.1.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
In this case, we show that w.h.p. AH achieves a success during the halving phase. We refer to a window with size in the range $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ as good. We first bound the latency required to reach a good window, and then analyze the expected collision cost.
Lemma 1
Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, AH terminates at or before a good window is reached.
Lemma 2
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, the latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
Proof
Under AH, each packet starts with an initial window size of $w_{0}=2^{\mathcal{C}^{\epsilon}}$. By assumption, the good window size is at most $w_{0}$, and so a good window will be reached during the halving phase. The halving phase executes, and by Lemma 1, w.h.p. in $\eta$ we are guaranteed to succeed at or before a good window. The number of windows that AH executes over before reaching a good window is at most $\lg(w_{0})=\mathcal{C}^{\epsilon}$, and the sample size for each such window is $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}}))\leq d\sqrt{\mathcal{C}}C^{\epsilon}$. Therefore, the latency is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
Lemma 3
(Biswas et al. [5]) Consider any slot $t$ where there exist $m$ active packets, each sending with the same probability $p$, where $p<1/m$. The probability of a collision in slot $t$ is at most $\frac{2m^{2}p^{2}}{\left(1-mp\right)}$.
Lemma 4
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, the expected collision cost for AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
Proof
In the halving phase, by Lemma 1, w.h.p. in $\eta$ a success occurs at or before a good window, which we are guaranteed by assumption to reach since $n\sqrt{\mathcal{C}}\leq w_{0}$. Thus, we consider any window of size $w\geq n\sqrt{\mathcal{C}}$ (and no more than $w_{0}$). We upper bound the probability of a collision using Lemma 3, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$).
Consider all the slots from the samples, and index these slots from $i\geq 1$. Let $X_{i}$ be an independent random variable where $X_{i}=1$ when the $i$-th slot contains a collision; $X_{i}=0$ otherwise. Then, by Lemma 3, we have:
$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}.$
Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ be the size of the first good window that is reached during halving phase. To get an upper bound on the expected collision cost, we first sum up the probabilities of a collision in any of the windows from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to $w^{*}\geq n\sqrt{\mathcal{C}}$, which requires at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$ halvings. Then, we multiply by the total number of slots in all the samples until a good window is reached, and then multiply by the cost of a collision.
To start, the sum of the probabilities is as follows:
$\sum_{j=0}^{C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2}\hskip-8.0pt\frac{2\left(\frac{n}{w_{0}/2^{j}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{j}}\right)\right)}$ $=\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\lg\mathcal{C})}{2}}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{j}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{j}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
$\leq\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\ln\mathcal{C})}{2}}\hskip 0.0pt\frac{2^{2j}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$
$\leq\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\ln\mathcal{C})}{2}}\hskip 0.0pt2^{2j}$
$=\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$
$<\frac{8}{\mathcal{C}}.$
We note that the second line by assumption of $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$. The third line holds for any $\mathcal{C}\geq 4$.
Recall that the sample size is $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=O(\mathcal{C}^{(1/2)+\epsilon})$. Over at most $\lg(w_{0})=O(C^{\epsilon})$ samples, the total number of slots in all the samples is $O(\mathcal{C}^{(1/2)+2\epsilon})$. Thus, letting $X=\sum_{i}X_{i}$ and using our upper bound on the probability above, the expected number of collisions is :
$E\left[X\right]$ $\leq E\left[\sum_{i=1}^{O(\mathcal{C}^{(1/2)+2\epsilon})}\hskip-8.0ptX_{i}\right]=O\left(\frac{\mathcal{C}^{(1/2)+2\epsilon}}{\mathcal{C}}\right)$
Thus, multiplying this by $\mathcal{C}$, we obtain that the expected collision cost is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
4.1.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
In this case, we show that w.h.p. in $\eta$ AH succeeds during the doubling phase. To begin, we show that $\mathcal{C}$ is small in $n$.
Lemma 5
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, $\mathcal{C}=O(\log^{1/\epsilon}n)$.
Proof
Since $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$, we have:
$\lg n$ $>\mathcal{C}^{\epsilon}-(\lg\mathcal{C})/2$
$>\mathcal{C}^{\epsilon}-(\mathcal{C}^{\epsilon}/2).$ (1)
To justify the second line, let $\delta\in(0,1)$, and note that:
$(\lg\mathcal{C})/2$ $<\mathcal{C}^{\epsilon}/2$
so long as $\epsilon>\frac{\lg\lg(\mathcal{C})}{\lg\mathcal{C}}$, which holds by assumption since $\epsilon$ is a constant, and this is well-defined since $\mathcal{C}\geq 4$. Returning to Equation 1, we have that:
$(2\lg n)^{1/\epsilon}$ $>\mathcal{C}.$
as desired, since $\epsilon$ is a constant.
Next, we show argue that a success is likely once we reach a window of size $\Theta(n)$.
Lemma 6
Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, in the doubling phase, AH achieves a success at or before a window of size $n$ is reached.
Lemma 7
Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, the latency required by AH is $O(\log^{4+(1/(2\epsilon))}n)$.
Proof
Under AH, all packets start with an initial window size of $w_{0}$, and they iteratively halve the window size until a window size of $2$ (in the worst case) is reached, or until a success, whichever occurs first. Pessimistically, assume no success occurs during the halving phase. By Lemma 2, the latency incurred for the halving phase is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
After that, the doubling phase begins. W.h.p. in $\eta$, by Lemma 6, a success at or before a window of size $n$; this requires $O(\log n)$ doublings and, for each window size, we have a sample of $O(\log(w_{\mbox{\tiny cur}}))=O(\log n)$ Therefore, the latency for the doubling phase is $O(\log^{2}n)$. It follows that the sum of the latency from the halving phase and the doubling phase is:
$O(\mathcal{C}^{(1/2)+2\epsilon}\log^{2}n)$ $=O(\log^{4+(1/(2\epsilon))}n)$
since $\mathcal{C}=O(\log^{1/\epsilon}n)$ by Lemma 5.
Lemma 8
Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, the collision cost is $O(\log^{4+(3/(2\epsilon))}n)$.
Proof
By assumption $n\sqrt{\mathcal{C}}>w_{0}$, so we can invoke Lemma 7, which states that w.h.p. in $\eta$ the latency is $O(\log^{4+(1/(2\epsilon))}n)$. Pessimistically, every slot may contain a collision, and by Lemma 5, $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, an upper bound on the total collision cost is:
$O(\mathcal{C}\,\log^{4+(1/(2\epsilon))}n))$ $=O(\log^{4+(3/(2\epsilon))}n)$
which yields the claim.
4.1.3 Total Cost of AH
We can now prove our upper bounds on the latency and expected collision cost for AH in the static setting.
Theorem 1.1. W.h.p. AH solves the static wakeup problem with the following costs:
  1. 1.
    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+2\epsilon}\right)$.
  2. 2.
    Else, the latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ latency and the expected collision cost is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
Proof
There are two cases to consider. In the first case, $n\sqrt{\mathcal{C}}\leq w_{0}$. By Lemmas 2 and 4, w.h.p. the latency and expected collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
In the second case, $n\sqrt{\mathcal{C}}>w_{0}$. By Lemmas, 7 and 8, w.h.p. when $n\sqrt{\mathcal{C}}>w_{0}$, AH has $O(\log^{4+(1/(2\epsilon))}n)$ latency and the collision cost is $O(\log^{4+(3/(2\epsilon))}n)$, yielding (2).
Finally, we note that the relationship between $n$ and $\mathcal{C}$— a relationship established by the adversary (recall Section 1.1)—dictates which upper bound applies. Since (1) dominates (2) when $C=\Omega(\log^{\Theta(1/\epsilon)}n)$, the result follows.
4.2 Lower Bound
[…]
Lemma 9
Any fair algorithm for the static wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
[…]
Lemma 10 (Biswas et al. [6], Lemma 24)
Fix any slot $t$ in any fair algorithm for the static wakeup problem. If $\texttt{Con}(t)>2$, then the probability of a collision is at least $1/10$.
By Lemma 10, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$. Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$. Under these circumstances, we provide a lemma that lower bounds the probability of a collision as a function of contention. We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
Lemma 11
Fix any slot $t$ where there are at least $2$ packets and $\texttt{Con}(t)\leq 2$ under the execution of any fair algorithm for the static wakeup problem. The probability of a collision in $t$ is $\Omega(\texttt{Con}^{2}(t))$.
Proof
For a fair algorithm, in any slot $t$, $\texttt{Con}(t)=np$. For any fixed slot, the probability of a collision involving exactly two packets (since there are at least $2$ packets by assumption) is a lower bound on the probability of a collision (that may involve more than $2$ packets), and this is $\binom{n}{2}p^{2}(1-p)^{n-2}$, which is at least:
$\geq\binom{n}{2}p^{2}(1-p)^{n}\geq\left(\frac{n^{2}}{2}\right)\left(\frac{\texttt{Con}(t)}{n}\right)^{2}e^{-2\cdot\texttt{Con}(t)}\geq\left(\frac{\texttt{Con}^{2}(t)}{2e^{2\cdot\texttt{Con}(t)}}\right)$
$=\Omega(\texttt{Con}^{2}(t))$
where the last line follows from $\texttt{Con}(t)\leq 2$.
We can derive a relationship between latency and expected collision cost for fair algorithms.
Theorem 1.2. Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$. Any fair algorithm that w.h.p. in $\eta$ solves the static wakeup problem with $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has w.h.p. an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-k\epsilon}/\log^{\beta}n)$.
Proof
Let $\mathcal{A}$ be an algorithm that w.h.p. in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is any constant. By Lemma 9, $\mathcal{A}$ has total contention:
$\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$ (2)
As we discussed, an implication of Lemma 10, we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot. Thus, we can use Lemma 11 (noting that $n\geq 2$), which implies that the expected collision cost in a single slot $t$ is:
[…]
for some constant $k>0$. So, the total expected collision cost is:
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\mathcal{C}$
By Jensen’s inequality for convex functions, we have:
$\frac{\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)}{L}\geq\left(\frac{\sum_{t=1}^{L}k\,\texttt{Con}(t)}{L}\right)^{2}$
which implies:
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\geq\frac{\left(\sum_{t=1}^{L}k\,\texttt{Con}(t)\right)^{2}}{L}$ (3)
So, by Equation 3, the expected cost is at least:
[…]
and since w.h.p., $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$, we have:
$\frac{\mathcal{C}\,k^{2}\,2^{2}}{L}$ $=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
which yields the claim.
Recalling Theorem 1.1, if the adversary selects $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then Theorem 1.2 implies that AH is within $O(\mathcal{C}^{4\epsilon})$-factor of the optimum expected collision cost for fair algorithms. For smaller values of $\mathcal{C}$, we are clearly within a polylogarithmic-in-n factor of the optimal latency and expected collision cost. These observations yield our Corollary 1.
4.3 Improved Bounds on Cost of Aim-High
AH achieves its bounds on latency and expected collision cost with high probability in $\eta$. However, note that if this guarantee fails, these quantities may be unbounded. This can happen in two ways. First, say that $n\sqrt{\mathcal{C}}\ll w_{0}$ and AH does not succeed in a good window. Then, the packets will complete the halving phase and execute the doubling phase, where the probability of a success can be vanishingly small. Second, when $n\sqrt{\mathcal{C}}>w_{0}$ and AH is unlucky and does not succeed within a window of size $\Theta(n)$, the packets may execute the doubling phase in perpetuity.
Intuitively, any fix requires periodically giving the packets another chance to succeed, for both cases: $n\sqrt{\mathcal{C}}\leq w_{0}$ and $n\sqrt{\mathcal{C}}>w_{0}$. Consequently, we modify AH in the following way.
Modifying AH. The new algorithm runs in iterations indexed by $j\geq 0$. For iteration $j=0$, the packets execute an entire halving phase. If there is no success, the packets then execute a truncated doubling phase; they run the original doubling phase for $2^{j}=1$ small samples (or until a success) starting at window size $4$; after this, iteration $j=0$ has ended. In the next iteration $j=1$, the packets execute another entire halving phase, and then execute another truncated doubling phase for $2^{j}=2$ small samples, again starting at window size $4$. And the process continues until the first success.
Intuition. If $n\sqrt{\mathcal{C}}\leq w_{0}$, then we now have a new chance of success in the halving phase of each iteration. Since success occurs w.h.p. in $\eta$, failing in the good window is such a low-probability event that the expected cost of the next $2^{j}$ small samples in the following truncated doubling phase is small.
Conversely, if $n\sqrt{\mathcal{C}}>w_{0}$, then the halving is useless, and we are searching for a window of size $\Theta(n)$ in the following truncated doubling phase of the iteration. Since $\mathcal{C}$ is $O(\log^{1/\epsilon}n)$ in this case, we can afford the extra latency and collision cost from the interleaved halving phases. Moreover, consider when an iteration is reached where the truncated doubling phase has packets increasing their window size to $\Theta(n)$; that is, once the truncated doubling phase runs for $\lg n$ samples (or, in other words, $j=\lg\lg n$). Then, the contribution to the expected cost from the next truncated doubling phase is made small by the low probability of failing in the current truncated doubling phase. The proofs for both statements below are in the appendix.
Lemma 12
Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency and expected collision cost are $O(\mathcal{C}^{(1/2)+2\epsilon})$.
Lemma 13
Assume $n\sqrt{\mathcal{C}}>w_{0}$. The expected latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
5 The Dynamic Setting
[…]
5.1 Technical Overview
Compared to the static setting, there are some significant differences in our analysis for the dynamic setting. For starters, there is no good window, given that multiple batches may be active, each with their respective windows. Instead, we have the analogous notion of good contention, which is contention in the range $3/\sqrt{\mathcal{C}},27/\sqrt{\mathcal{C}}]$. In Lemma 15, we show that a success is likely if we have $\Omega(\sqrt{\mathcal{C}})$ slots with good contention. At time slot $t$, contention is said to high when $\texttt{Con}(t)>27/\sqrt{\mathcal{C}}$, and low when $\texttt{Con}(t)<3/\sqrt{\mathcal{C}}$.
Unlike the static setting, we cannot divide our analysis into two cases (i.e., $n\sqrt{\mathcal{C}}\leq w_{0}$ and $n\sqrt{\mathcal{C}}>w_{0}$) quite as cleanly. For instance, in the dynamic setting, suppose that $n\sqrt{\mathcal{C}}>w_{0}$ and the adversary initially injects a small number of packets $n^{\prime}\leq w_{0}/\sqrt{\mathcal{C}}$ which execute the halving phase.
  • $\bullet$
    If no more packets are activated, these $n^{\prime}$ packets may reach good contention and yield a success. This is despite $n\sqrt{\mathcal{C}}>w_{0}$, which would not have occurred in the static setting.
  • $\bullet$
    Else, suppose the adversary injects the remaining packets just prior to these $n^{\prime}$ packets reaching good contention. All of sudden, the contention jumps spikes (preventing a success) in a manner that was not possible in the static setting.
[…]
Given that the sending probabilities of packets in separate batches will differ, a technical challenge arises when trying to bound the probability of a collision. We tackle this via an exchange argument in the proof of Lemma 16, and then in Lemma 17 we leverage this to bound the collision cost as a function of contention.
A similar challenge is addressed in trying to show that good contention can be reached during the halving phase for $t^{*}$ sufficiently large. Again, the dynamic aspect of the problem prevents a straightforward analysis. Thus, we argue via Lemma 18 that, if there had been high contention in the halving phase, we would been able to “rewind” backwards in the execution and show that there was a period of good contention leading to a success and termination; in other words, we can derive a contradiction.
Error. In the static case, a success occurs when all $n$ packets have a window large enough such that using $\ell=\ln(w_{\mbox{\tiny cur}})$ in the sample size yields guarantees that are w.h.p. in $\eta$. However, in the dynamic setting, this is not the case; a success may occur when the currently-active packets (in an assortment of batch sizes) have small respective windows. Thus, in this setting, AH succeeds with probability at least $1-O(\mathcal{C}^{-d^{\prime}})$, where $d^{\prime}>0$ is a constant that can be made arbitrarily large depending only on sufficiently large $d$.
5.2 Upper Bound
Throughout, we index slots starting at $1$, which corresponds to the slot where the first packet is injected; that is, we ignore all possible empty slots prior to the first packet.
Since the minimum window size is $2$ under the execution of AH, the largest sending probability is $1/2$, and so we make use of the following prior result for both cases $t\leq t^{*}$ and $t>t^{*}$.
Lemma 14
(Bender et al [4]) Assume for any fixed slot $t$ that the sending probability of each active packet is at most $1/2$. Then, the probability of a success in slot $t$ is at least $\texttt{Con}(t)/e^{2\texttt{Con}(t)}$.
5.2.1 Case 1: $t\leq t^{*}$.
To show w.h.p. in $\mathcal{C}$ a success occurs, we will need the following analog to Lemma 1.
Lemma 15
After at least $d\sqrt{\mathcal{C}}\ln\mathcal{C}$ slots with good contention, w.h.p. in $\mathcal{C}$ a packet has succeeded.
Our analysis of collision cost for this case depends on the following. For $n$ probability values $p_{1},p_{2},...,p_{n}$, consider the following:
$S=\sum_{i\neq j\neq\cdots\neq k}\overbrace{p_{i}p_{j}\dots p_{k}}^{\alpha}$
which is the sum of all products of $\alpha$ probability values for $1\leq\alpha\leq n$. The next lemma addresses the maximum value of this sum.
Lemma 16
There is a setting of the $n$ probability values to be equal that maximizes $S$.
Proof
Consider any setting of the probability values that maximizes $S$. If this setting already has all the probability values set equal, then we are done. Else, there exist two indices $i$ and $j$ such that $p_{i}<p_{j}$. We now present an exchange argument.
Consider a small shift of $\delta=\frac{p_{j}-p_{i}}{2}$ from $p_{j}$ to $p_{i}$, keeping all other probability values the same. That is, let ${p_{i}}\leftarrow p_{i}+\delta$ and ${p_{j}}\leftarrow p_{j}-\delta$. This redistribution affects any terms in $S$ that involve $p_{i}$ or $p_{j}$. Consequently, we analyze its impact by considering two cases: terms that involve both $p_{i}$ and $p_{j}$, and terms that involve only one of them.
At first, we consider the change in $S$ for the first case. Each summand in $S$ involving both $p_{i}$ and $p_{j}$ changes as follows:
$(p_{i}+\delta)(p_{j}-\delta)\cdots p_{k}=(p_{i}p_{j}-\delta p_{i}+\delta p_{j}-\delta^{2})\cdots p_{k}.$
Focusing on the first factor on the right, we have:
 $\mbox{\hskip 8.0pt}(p_{i}p_{j}-\delta p_{i}+\delta p_{j}-\delta^{2})$
$=p_{i}p_{j}-\left(\frac{p_{j}-p_{i}}{2}\right)p_{i}+\left(\frac{p_{j}-p_{i}}{2}\right)p_{j}-\left(\frac{p_{j}-p_{i}}{2}\right)^{2}$
$=\frac{1}{4}\left(4p_{i}p_{j}-2p_{i}p_{j}+2p^{2}_{i}+2p^{2}_{j}-2p_{i}p_{j}-p^{2}_{j}-p^{2}_{i}+2p_{i}p_{j}\right)$
$=\frac{1}{4}\left(p^{2}_{j}+p^{2}_{i}+2p_{i}p_{j}\right)$
where the last line is strictly positive.
Now, for the second case, note that for every summand with $p_{i}$ and not $p_{j}$, there is a “matching” term with $p_{j}$ and not $p_{i}$, and with all other probability terms in the summand the same. Thus, we analyze the increase due to $(p_{i}+\delta)p_{l}\cdots p_{k}$ and the decrease from the matching term $(p_{j}-\delta)p_{l}\cdots p_{k}$, and show that the net result is at least 0. Formally, is the following true?
$\left((p_{i}+\delta)p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)\hskip-1.0pt-\hskip-1.0pt\left(p_{i}p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)\hskip-1.0pt\geq\hskip-1.0pt\left(p_{j}p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)\hskip-1.0pt-\hskip-1.0pt\left((p_{j}-\delta)p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)$
Observe that by canceling out a factor of $p_{l}\cdots p_{k}$, the inequality is true. Thus, by the above analysis, the net change is positive, which contradicts our assumption that $S$ was already maximized.
Packets in different batches have different sending probabilities, which complicates our analysis. To this end, the next lemma captures the probability of a collision as a function of contention.
Lemma 17
For any slot with low or good contention, the probability of a collision is at most $O(1/\mathcal{C})$.
Proof
We establish this argument by comparing collisions involving different numbers of packets. Collisions may occur between pairs, triples, quadruples, and so on. The probability of an $\alpha$-way collision—a collision involving any $\alpha\geq 2$ packets—is at most
$\leq\sum_{i\neq j\neq\dots\neq k}\overbrace{p_{i}p_{j}\cdots p_{k}}^{\alpha}$
$\leq\sum_{1}^{\binom{n}{\alpha}}(p^{*})^{\alpha}$
$=\binom{n}{\alpha}\left(\frac{27}{n\sqrt{\mathcal{C}}}\right)^{\alpha}$
$=O\left(\frac{1}{\mathcal{C}^{\alpha/2}}\right)$ (4)
The second line follows from Lemma 16, where $p^{*}$ denotes the common sending probability of all packets. The third line follows from the definition of the upper bound of good $\texttt{Con}(t)=27/\sqrt{\mathcal{C}}$ (which, of course, is an upper bound on low contention).
In any fixed slot, we compute the probability of collisions involving exactly two packets by setting $\alpha=2$ in Equation 4, which yields a bound of $O(1/\mathcal{C})$. Similarly, the total probability of collision involving three or more packets ($\alpha\geq 3$) is:
$O\left(\sum_{\alpha=0}^{n}\left(\frac{1}{\sqrt{\mathcal{C}}}\right)^{\alpha}\right)$ $=O\left(\sum_{\alpha=3}^{n}\left(\frac{1}{\sqrt{\mathcal{C}}}\right)^{\alpha}-1-\frac{1}{\sqrt{\mathcal{C}}}-\frac{1}{\mathcal{C}}\right)$
$=O\left(\frac{1}{\mathcal{C}^{3/2}}\right)$
Therefore, the total probability of collisions at most $\Theta(1/\mathcal{C})+O(1/(\mathcal{C}^{3/2}))=O(1/\mathcal{C})$.
Recall that for $t\leq t^{*}$, $n(t)\leq w_{0}/\sqrt{\mathcal{C}}$ and, here, we wish to argue that a success will occur while packets are executing their respective halving phases. Note that the contention can be good or low in the initial window; pessimistically, consider the case where it is low. Moreover, at some point in the halving process, as the windows shrink in size, contention will be high. We need to argue that, in between these points of low and high contention, there is a period of good contention. Our next lemma provides a tool for showing this is true; this is done by considering a point in time in the halving phase where contention can be high (or good) and “rewinding” backwards in the execution, while tracking contention. Let s $d\sqrt{\mathcal{C}}\ell=d\sqrt{\mathcal{C}}\ln\mathcal{C}$ denote the large sample size used during the halving phase of in AH (recall Section 3).
Lemma 18
Suppose at time $t\leq t^{*}$, all active packets are executing the halving phase, and $\texttt{Con}(t)\geq 3/\sqrt{\mathcal{C}}$. Then:
$C(t-s)\in\left[\frac{\texttt{Con}(t)}{3},\frac{\texttt{Con}(t)}{2}\right].$
Proof
Rewinding by $s$ slots from the current time $t$, each batch that exists at time $t$ either (1) still exists back at time tst-s, but with windows twice as large, or (2) does not exist because it has not yet been injected back at time tst-s. In the case of (1), since all batches exist back at time tst-s with their respective windows doubled, then $C(t-s)\leq\texttt{Con}(t)/2$.
Else, in the case of (2), some packets that are present at time $t$ are not yet injected at time tst-s. To measure the change in contention, we ask: how much contention is present at time $t$, but absent back at time tst-s due to these packets not yet active? Note that such packets can only have been in the system for at most $s$ slots (otherwise, they would still be present at time tst-s); therefore, each of these packets has a window size of $w_{0}$. Since by assumption $n(t)\leq w_{0}/\sqrt{\mathcal{C}}$, this means that the contention contributed by all such packets at time $t$ is at most $n(t)/w_{0}\leq 1/\sqrt{\mathcal{C}}$. Thus, we can get a lower bound on the contention at time tst-s by first subtracting out this contribution, and then accounting for the doubling of the windows for those packets that do exist back at time tst-s. That is, $C(t-s)\geq\frac{\texttt{Con}(t)-(1/\sqrt{\mathcal{C}})}{2}$. To simplif this , we note that $\frac{\texttt{Con}(t)-(1/\sqrt{\mathcal{C}})}{2}$ $\geq\texttt{Con}(t)/3$, since by assumption $\texttt{Con}(t)\geq 3/\sqrt{\mathcal{C}}$.
[…]
Lemma 19
Suppose that $t^{*}>k\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C}$ for a sufficienty large constant $k>0$. Then, w.h.p. in $\mathcal{C}$, AH yields a success in $\mathcal{I}$ with latency and expected collision cost $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
5.2.2 Case 2: $t>t^{*}$.
We start with an analog to Lemma 5.
Lemma 20
Assume $t>t^{*}$. Then, $\mathcal{C}=O(\log^{1/\epsilon}n)$.
Proof
By definition of $t^{*}$, if $t>t^{*}$, then $n(t)>w_{0}/\sqrt{\mathcal{C}}$. Solving for $\mathcal{C}$ is done as in the proof of Lemma 5 yields $\mathcal{C}=O(\log^{1/\epsilon}n(t))$, and the result follows from noting that $n(t)\leq n$ for any slot $t$.
We also need an analog to Lemma 6 (used in the static setting), which shows that $d\ell$ slots with constant contention (rather than a window of size $n$) yields a success with high probability in $\ell$. For this, we define the interval of adequate contention to be $[1/16,1]$.
Lemma 21
Over $d\ln\mathcal{C}$ slots of adequate contention, w.h.p. in $\mathcal{C}$, a success occurs.
For a sufficiently large constant $\kappa>0$, define a gap to be a contiguous set of at least $\gamma$ = $\kappa(\log^{1+(1/(2\epsilon))}n)\log\log n$ slots, during which the adversary makes no packet injections. We note that the value of $\kappa$ is lower bounded in the proof of Lemma 22 below.
Lemma 22
Consider any time slot $t>t^{*}$ that marks the start of a gap. Then, w.h.p. in $\mathcal{C}$, a success occurs during the gap.
Lemma 23
W.h.p. in $\mathcal{C}$, a success occurs within $O(n(\log^{1+(1/(2\epsilon))}n)$ $\log\log n)$ slots of $t^{*}$ and with an associated $O(n(\log^{1+(3/(2\epsilon))}n)$ $\log\log n)$ collision cost over these slots.
Proof
Given a gap starting in any slot $t>t^{*}$, by Lemma 22 w.h.p. in $\mathcal{C}$, there is a success in this gap. Pessimistically, assume no success occurs before the first gap (otherwise, our latency and collision cost is smaller). Starting from slot $t^{*}+1$, what is the maximum number of slots that the adversary can force to elapse before the first gap must occur?
To avoid a gap from occurring, a minimum of one packet must be activated at least every $\gamma-1$ slots. Given that there are at most $n$ packets ever activated, this implies that the number of slots prior to the first gap and a consequent success is at most $n\gamma$ $O(n(\log^{1+(1/(2\epsilon))}n)\log\log n)$ slots.
The collision cost is upper bounded by considering a collision in every slot, which is:
$O(\mathcal{C}n(\log^{1+(1/(2\epsilon))}n)\log\log n)$ $=O(n(\log^{1+(3/(2\epsilon))}n)\log\log n)$
by Lemma 20, which completes the argument.
5.2.3 Total Cost of AH
We can now prove bounds on latency and the expected collision cost of AH for the dynamic setting.
Theorem 1.3. W.h.p. in $\mathcal{C}$, AH solves the dynamic wakeup problem with the following costs:
  1. 1.
    If  $\mathcal{C}=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$ for a small constant $\delta>0$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+\epsilon}\log n\right)$,
  2. 2.
    Else, the latency is $O(n(\log^{1+(1/(2\epsilon))}n)\log\log n)$ and the expected collision cost is $O(n(\log^{1+(3/(2\epsilon))}n)\log\log n)$.
Proof
If $t^{*}>k\mathcal{C}^{(1/2)+\epsilon}\log\mathcal{C}$, for a sufficiently large constant $k>0$, Lemma 19 implies that w.h.p. in $\mathcal{C}$, a success occurs with latency and expected collision cost $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
Else, by Lemma 23, w.h.p. in $\mathcal{C}$, a success occurs within $O(n$ $(\log^{1+(1/(2\epsilon))}n)$ $\log\log n)$ slots of $t^{*}$ and with an associated $O(n$ $(\log^{1+(3/(2\epsilon))}n)$ $\log\log n)$ collision cost over these slots. What about the additional latency and collision cost up to slot $t^{*}$? By Lemmas 19 and 20, each of these is:
$O(\mathcal{C}^{(1/2)+\epsilon}\log\mathcal{C})=O((\log^{1+(1/(2\epsilon))}n)\log\log n)$
which is asymptotically smaller, and so the claimed bounds follow.
Finally, given as input a (large) value of $n$ to the algorithm, the adversary may choose which upper bound applies for any given execution by setting $\mathcal{C}$ and $t^{*}$. For sufficiently large $\mathcal{C}$, the bounds in (1) dominate the bounds in (2). Specifically, this occurs for:
$\mathcal{C}$ $=\Omega(n^{1/(0.5+\epsilon)}\log^{\Theta(1/\epsilon))}n)=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$
where the last line follows for some small constant $\delta>0$.
We note that the costs stated in (1) of Theorem 1.3 differ from those in the corresponding (1) of Theorem 1.1. This occurs because of the different values of $\ell$ (recall Section 3). In the dynamic case, $\ell=\ln\mathcal{C}$, which is used to get our w.h.p. in $\mathcal{C}$ guarantees, and is the source of the $O(\log n)$-factor. In contrast, for the static case $\ell=\ln(w_{\mbox{\tiny cur}})$, which is used to get our w.h.p. in $\eta$ guarantees, and is the source of the additional $O(\mathcal{C}^{\epsilon})$-factor.
5.3 Lower Bound
In the dynamic setting, Jurdziński and Stachowiak [18] highlight the class of algorithms where all packets exhibit identical behavior, governed by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability in the $i$-th time slot after a packet is activated; we refer to this class of algorithms as batch-fair.
Observe that, when restricted to the case where all packets start in slot $1$ (i.e., the static setting), a batch-fair algorithm is fair. Thus, we can apply Theorem 1.2 here to obtain the following.
Lemma 24
Any batch-fair algorithm that w.h.p. solves dynamic wakeup with $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$ latency must incur an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-\epsilon}/\log n)$.
Note that AH is a batch-fair algorithm for the dynamic case. By Theorem 1.3, if the adversary selects $C$ sufficiently large—that is, $\mathcal{C}=\Omega(n^{2-\delta})$—then Lemma 24 implies that AH is within an $O(\mathcal{C}^{2\epsilon}\log^{2}n)$-factor of the optimal expected collision cost.
For smaller values of $\mathcal{C}$, however, Lemma 24 is not useful. In this case, we consider a prior result by Jurdziński and Stachowiak [18], who show a lower bound on latency of $\Omega(n/\log n)$ for batch-fair algorithms, which applies to our setting. Therefore, for smaller values of $\mathcal{C}$, AH is within an $O(\log^{\Theta(1/\epsilon)}n)$-factor of the optimal latency. These observations yield Theorem 1.4.
6 Future Work
For the dynamic setting, an interesting open problem is whether we can achieve finite expected latency and collision cost bounds that hold with probability $1$. We would also like to prove general lower bounds for the static and dynamic settings. Finally, it may be worthwhile to explore resilience to jamming under this model.
Appendix
Here, we provide arguments that could not be included in the main paper due to space constraints.
Proofs from Section 4.1
Lemma 1. Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, AH terminates at or before a good window is reached.
Proof
If a success occurs before a good window, then the claim is true. Else, we are at a good window with size $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3), which occurs within $O(\log w_{0})$ halvings. Here, the probability of a success in any fixed slot is at least:
[…]
$\geq\left(\frac{1}{3e\sqrt{\mathcal{C}}}\right)\mbox{~~~since $\mathcal{C}\geq 4$.}$
Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$. The expected number of successes in the sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots is at least:
$E[X]\geq\left(\frac{d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})}{3e\sqrt{\mathcal{C}}}\right)$ $=\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$
We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
$Pr\left(X\leq(1-\delta)\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$ $\leq Pr\left(X\leq\frac{d\ln(w_{\mbox{\tiny cur}})}{6e}\right)$
$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
$\leq\eta^{-d^{\prime\prime}}$
where the third line follows from $w_{\mbox{\tiny cur}}=\Theta(\ln n\sqrt{\mathcal{C}})$, and $d^{\prime}$ and $d^{\prime\prime}$ can be arbitrarily large constants for $d$ is sufficiently large.
Lemma 6. Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, in the doubling phase, AH achieves a success at or before a window of size $n$ is reached.
Proof
Suppose we are in a window of size $w=n$. The probability of a success in any fixed slot is at least:
[…]
$\geq 1/e^{2}.$
Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$. The expected number of successes in the sample over $d\ln(w_{\mbox{\tiny cur}})$ is at least:
$E[X]\geq\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{e^{2}}\right).$
We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
$Pr\left(X\leq(1-(1/2))\frac{d\ln(w_{\mbox{\tiny cur}})}{e^{2}}\right)$ $\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)e^{2}}}$
$=e^{-\frac{d\ln(n)}{10e^{2}}}$
$\leq n^{-d^{\prime}}$
$\leq\eta^{-d^{\prime\prime}}$
The second follows from $w_{\mbox{\tiny cur}}=\Theta(n)$, and where $d^{\prime\prime}$ and $d^{\prime}$ can be arbitrarily large constant depending only on $d$ being sufficiently large. The last line follows from Lemma 5, which implies that $n\gg\mathcal{C}$.
Proofs from Section 4.2
Lemma 9. Any fair algorithm for the static wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
Proof
[…]
$TC(\mathcal{A})=\sum_{u}LC(u)=n\cdot LC(u)$
By our assumption, $n\cdot LC(u)<2$. The probability of $u$ not succeeding in slot any fixed $t$ is at least 1p(t)1-p(t). Therefore, the probability that $u$ never succeeds over the execution of $\mathcal{A}$ is:
$\geq\prod_{i=1}^{\lambda}(1-p(i))$
$\geq e^{-2\sum_{i}^{\lambda}p(i)}$
$\geq e^{-2\cdot LC(u)}$
$\geq e^{-2\cdot TC(\mathcal{A})/n}$
$>e^{-4/n}$
where the second line follows from Fact 1(b), and the third line follows from the definition of $LC(u)$. The last line follows from the assumption that $TC(\mathcal{A})<2$.
It follows that the probability of all packets failing over the execution of $\mathcal{A}$ is:
$\geq\left(e^{-4/n}\right)^{n}=\frac{1}{e^{4}}$
Thus, there is a constant probability of all packets failing, which contradicts the w.h.p. guarantee of $\mathcal{A}$.
Proofs from Section 4.3
Lemma 12. Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency and expected collision cost are $O(\mathcal{C}^{(1/2)+2\epsilon})$.
Proof
In the halving phase of iteration $j$, the latency and collision cost is $O(\mathcal{C}^{(1/2)+2\epsilon})$ by Lemmas 2 and 4.
In the truncated doubling phase of iteration $j$, the latency is
$O(2^{j}\ell)$ $=O(2^{j}\log(w_{\mbox{\tiny cur}}))$
$=O(2^{j}\log(2^{j}))$
$=O(j2^{j})$
and the corresponding collision cost is at most $O(j2^{j}\mathcal{C})$, since in the worst case there is a collision in every slot; note that this dominates the latency.
By Lemma 1, for any fixed halving phase, the probability of failing in a good window is $O(1/{\texttt{poly}({\eta})})$, where we can make the degree of the polynomial an arbitrarily large constant depending on sufficiently large $d$. Thus, each of the expected latency and collision cost is at most: the bound from Case (1) of Theorem 1.1 for the w.h.p. in $\eta$ event of succeeding in iteration $j=0$, plus the increasingly low-probability event of failing up until the $j$-th iteration for $j\geq 1$ times the associated cost. That is:
$=O\left(\mathcal{C}^{(1/2)+2\epsilon}\right)+\sum_{j=1}^{\infty}O\left(\frac{1}{{\texttt{poly}({\eta})}}\right)^{j}O\left(j2^{j}\mathcal{C}\right)$
$=O\left(\mathcal{C}^{(1/2)+2\epsilon}\right)$
which completes the argument.
Lemma 13. Assume $n\sqrt{\mathcal{C}}>w_{0}$. The expected latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
Proof
The latency and collision cost for the halving phase of iteration $j$ are each $O(\mathcal{C}^{(1/2)+2\epsilon})$ by Lemmas 2 and 4, and this becomes $O(\log^{O(1/\epsilon)}n)$ by Lemma 5.
For the truncated doubling phase of iteration $j$, the latency is
$O(2^{j}\ell)$ $=O(2^{j}\log(w_{\mbox{\tiny cur}}))$
$=O(2^{j}\log(2^{j}))$
$=O(j2^{j})$
The corresponding collision cost is at most $O(j2^{j}\mathcal{C})=O(j2^{j}\log^{1/\epsilon}n)$, since in the worst case there is a collision in every slot, and we can invoke Lemma 5; note that this dominates the latency.
A window of size $n$ is reached by at most the $\lg\lg(n)$-th iteration; $2^{j}$ samples are executed, each sample doubles the window, and we need only $\lg(n)$ doublings. Thus, the latency and collision cost up to a window of size $n$ is $O((\log^{1/\epsilon}n){\texttt{poly}({\log n})})=O(\log^{\Theta(1/\epsilon)}n)$.
By Lemma 6, in a window of size $n$, the probability of failure is $O(1/{\texttt{poly}({\eta})})$. Using Case (2) of Theorem 1.1, the expected cost is this:
$=O(\log^{\Theta\left(\frac{1}{\epsilon}\right)}n)+\hskip-5.0pt\sum_{j=\lg\lg n}^{\infty}O\left(\frac{1}{{\texttt{poly}({\eta})}}\right)^{j-\lg\lg(n)+1}O\left(j2^{j}\log^{\frac{1}{\epsilon}}n\right)$
$=O(\log^{\Theta(1/\epsilon)}n)$
which completes the argument.
Proofs From Section 5.2
Lemma 15. After at least $d\sqrt{\mathcal{C}}\ln\mathcal{C}$ slots with good contention, w.h.p. in $\mathcal{C}$ a packet has succeeded.
Proof
By the definition of good contention, $3/\sqrt{\mathcal{C}}\leq\texttt{Con}(t)\leq 27/\sqrt{\mathcal{C}}$. Thus, by Lemma 14, we have that the probability of a success in slot $t$ is:
$\frac{\texttt{Con}(t)}{e^{2\texttt{Con}(t)}}$ $=\frac{b}{\sqrt{\mathcal{C}}}.$
for some constant $b>0$. The probability of failure over $d\sqrt{\mathcal{C}}\ln{n}$ slots is:
$\left(1-\frac{b}{\sqrt{\mathcal{C}}}\right)^{d\sqrt{\mathcal{C}}\ln{\mathcal{C}}}$ $\leq e^{-d^{\prime}\ln{\mathcal{C}}}$
$\leq\frac{1}{\mathcal{C}^{d^{\prime}}}$
where the first line follows from Fact 1 (a), and the constant $d^{\prime}$ is can be made arbitrarily large depending on sufficiently large $d$.
Lemma 22. Consider any time slot $t>t^{*}$ that marks the start of a gap. Then, w.h.p. in $\mathcal{C}$, a success occurs during the gap.
Proof
Let $B$ be the most-recent batch injected prior to the gap starting at $t$. At some later slot $t^{\prime}$, the packets in $B$ will have completed the halving phase; this requires a number of slots at most:
 $\mbox{\hskip 10.0pt}s\lg(w_{0})$
$=(d\sqrt{\mathcal{C}}\ln\mathcal{C})\lg(w_{0})$
$=O((\log^{1/(2\epsilon)}n)(\log\log n)\mathcal{C}^{\epsilon})$
$=O((\log^{1/(2\epsilon)}n)(\log\log n)(\log n))$
$=O((\log^{1+(1/(2\epsilon))}n)\log\log n)$
using Lemma 20.
Since $B$ was the last batch injected prior to the gap starting, we know at slot $t^{\prime}$ that all active packets are executing the doubling phase. We also know that contention is high at slot $t^{\prime}$, since any packet in $B$ has a window of size $4$ (at the start of the doubling phase) and, thus, contributes at least $1/2$ to contention.
During the doubling phase for any fixed batch, over each small sample of $\sigma=d\ln\mathcal{C}$ slots, the corresponding sending probability for those packets is halved. Due to different injection times, these small samples may be unsynchronized (recall Section 3.1.2). However, in every $s^{\prime}=2\sigma$ contiguous slots, all batches are guaranteed to halve their sending probability at least once and at most twice. Thus, starting at slot $t$, after every $s^{\prime}$ slots, the contention from packets in the system is reduced by a constant factor between $2$ and $4$.
Given that adequate contention spans a factor of $16$ (recall, $[1/16,1]$), the contention will enter this range (i.e., it cannot be skipped), and this occurs within at most:
$s^{\prime}O(\lg(\texttt{Con}(t)))$ $\leq O(s^{\prime}\lg(n))$
$=O(\log n\log\log n)$
slots, where the second line follows given that contention is always upper bounded by $n$ and Lemma 20.
Additionally, again, given that the range of adequate contention spans a factor of $16$, at least one full small sample will execute with good contention; that is, contention will remain good for at least $d\ln\mathcal{C}$ slots. Therefore, by Lemma 21, w.h.p. in $\mathcal{C}$ there is a success.
The total latency measured from slot $t$ is at most:
$=O((\log^{1+(1/(2\epsilon))}n)\log\log n)+O(\log n\log\log n)$
$\leq k(\log^{1+(1/(2\epsilon))}n)\log\log n$
where $k$ is a sufficiently large positive constant. Thus, for $\kappa$ (in the definition of a gap) chosen such that $\kappa\geq k$, the claim holds.
Lemma 19. Suppose that $t^{*}>k\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C}$ for a sufficienty large constant $k>0$. Then, w.h.p. in $\mathcal{C}$, AH yields a success in $\mathcal{I}$ with latency and expected collision cost $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
Proof
Since $n(t)\leq w_{0}/\sqrt{\mathcal{C}}$ for $t\leq t^{*}$, the contention in slot $1$ is $n(1)/w_{0}\leq 1/\sqrt{\mathcal{C}}$. Thus, at the start of the execution, contention can only be low or good (but not high).
Next, we argue that within $\mathcal{I}$ the contention cannot change from low to high without good contention slots in between. Over any two consecutive slots in $\mathcal{I}$, the contention from active packets may double if all the corresponding windows halve. Also, again, the largest possible increase in contention due to newly-injected packets is $n(t)/w_{0}$ $\leq 1/\sqrt{\mathcal{C}}$, since $t\leq t^{*}$. Thus, given that good contention spans a factor of $9$ (recall $3/\sqrt{\mathcal{C}},27/\sqrt{\mathcal{C}}]$), contention cannot transition from low to high without at least some period of good contention in between.
Do we have high contention in $\mathcal{I}$? Consider the first batch, $B$, injected in slot $1$. Barring a success, the packets in $B$ will execute their halving phase until a window of size $2$ is reached within $t^{\prime\prime}=O(s\log w_{0})=O(C^{(1/2)+\epsilon}\log\mathcal{C})$ slots, at which point they contribute high contention. By the above, this implies a period of good contention in $\mathcal{I}$.
How many slots of good contention are in $\mathcal{I}$? There are two possibilities, but we will show that only the first can occur. In Possibility (1), at least $s=d\sqrt{\mathcal{C}}\ln\mathcal{C}$ slots of good contention occur before any slots with high contention. In Possibility (2), less than $s$ slots of good contention occur before any slots with high contention.
For Possibility (1), Lemma 15 guarantees w.h.p. in $\mathcal{C}$ a success by the end of those $s$ slots of good contention and, as derived above, this occurs by slot $t^{\prime\prime}$. Thus, the latency is $O(\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C})$ $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$ given that $C={\texttt{poly}({n})}$ (recall Section 1.1). By Lemma 17, since contention up to this point is low or good, the probability of a collision up until this termination is $O(1/C)$, yielding $O((1/\mathcal{C})\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C})$ collisions in expectation, and thus an expected collision cost of $O(\mathcal{C}^{(1/2)+\epsilon}\log\mathcal{C})$ $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
For Possibility (2), let $t^{\prime}\in\mathcal{I}$ be the first slot with high contention; this occurs before packets in $B$ finish their halving phase (since, again, $B$ contributes high contention when window size $2$ is reached). Since $B$ is the first batch injected, at $t^{\prime}$ all active packets are executing a halving phase. Thus, we can use Lemma 18.
Starting at $t^{\prime}$, while contention remains at least $3/\sqrt{\mathcal{C}}$, for every $s$ slots we rewind into the past, the contention decreases by at least a factor of $2$ and at most a factor of $3$. We repeatedly invoke Lemma 18 in order to rewind to a point in time where contention is in $3/\sqrt{\mathcal{C}},9/\sqrt{\mathcal{C}}]$, which we claim occurs in $\mathcal{I}$. To see this, note that:
$\texttt{Con}(t^{\prime})$ $\leq n(t^{\prime})/2\leq w_{0}/(2\sqrt{\mathcal{C}})$
since $t^{\prime}\in\mathcal{I}$. Thus, dividing the maximum possible contention by $2$ at most $\mathcal{C}^{\epsilon}$ times lands us in $3/\sqrt{\mathcal{C}},9/\sqrt{\mathcal{C}}]$. That is, we must rewind by at most $\mathcal{C}^{\epsilon}$ samples – at most $dC^{(1/2)+\epsilon}\ln\mathcal{C}$ slots – which is accommodated by $t^{*}>kC^{(1/2)+\epsilon}\ln\mathcal{C}$, for $k$ sufficiently large.
Finally, even if batches have unsynchronized samples (recall Section 3.1.2), we are guaranteed at least $s$ slots with contention in the range $3/\sqrt{\mathcal{C}},27/\sqrt{\mathcal{C}}]$. Thus, at least $s$ slots of good contention did occur within $\mathcal{I}$ prior to high contention, which contradicts Possibility (2).
Lemma 21. Over $d\ln\mathcal{C}$ slots of adequate contention, w.h.p. in $\mathcal{C}$, a success occurs.
Proof
Using Lemma 14, the probability of success in any slot $t$ with adequate contention is $\texttt{Con}(t)/e^{2\texttt{Con}(t)}\geq 1/(16e^{1/8})>1/20$. The probability of no success over all $d\ln\mathcal{C}$ slots is at most:
$(1-1/20)^{d\ln\mathcal{C}}$ $\leq e^{d\mathcal{C}/20}$
$=\mathcal{C}^{-d^{\prime}}$
where the first line follows by Fact 1(a), and where the constant $d^{\prime}>0$ can be made arbitrarily large for a sufficiently large $d$.

### v2

@@
-A Gentle Wakeup Call: Symmetry Breaking with Less Collision Cost
+A Gentle Wakeup Call Symmetry Breaking with Less Collision Cost
 Abstract
 The wakeup problem addresses the fundamental challenge of symmetry breaking.
-There are $n$ devices sharing a time-slotted multiple access channel.
-In any fixed slot, if a single device sends a packet, it succeeds; however, if two or more devices send, then there is a collision and none of the corresponding packets succeed.
-For the static version of wakeup, all packets are initially active (i.e., can send and listen on the channel); for the dynamic version, the packets become active at arbitrary times.
-In both versions, the goal is to successfully send a single packet.
-Prior results on wakeup have largely focused on the number of slots until the first success; that is, the latency.
-However, in many modern systems, collisions introduce significant delay, an aspect that current wakeup algorithms do not address.
-For instance, while existing results for static wakeup have polylogarithmic-in-$n$ latency, they can incur additional latency that is linear in the cost of a collision $\mathcal{C}$.
-Thus, the total latency is large and dominated by the contributions from collisions.
+Initially, $n$ devices share a time-slotted multiple access channel, which models wireless communication.
+A transmission succeeds if exactly one device sends in a slot; if two or more transmit, a collision occurs and none succeed.
+The goal is to achieve a single successful transmission efficiently.
+Prior work on wakeup primarily analyzes latency—the number of slots until the first success.
+However, in many modern systems, each collision incurs a nontrivial delay, $\mathcal{C}$, which prior analyses neglect.
+Consequently, although existing algorithms achieve polylogarithmic-in-$n$ latency, they still suffer a delay of $\Omega(\mathcal{C})$ due to collisions.
 Here, we design and analyze a randomized wakeup algorithm, Aim-High.
-For sufficiently large $\mathcal{C}$ and with bounded error, Aim-High has latency and expected collision cost that is nearly $O(\sqrt{\mathcal{C}})$ for both the static and dynamic versions.
-Otherwise, the latency and expected collision cost are $O({\texttt{poly}({\log n})})$ for the static setting, and $O(n\,{\texttt{poly}({\log n})})$ for the dynamic setting.
-We also establish lower bounds that complement these results.
-Keywords:
-Wakeup, symmetry breaking, multiple access channel, collision cost
+When $\mathcal{C}$ is sufficiently large with respect to $n$, Aim-High has expected latency and expected total cost of collisions that are nearly $O(\sqrt{\mathcal{C}})$; otherwise, both quantities are $O({\texttt{poly}({\log n})})$.
+Finally, for a well-studied class of algorithms, we establish a trade-off between latency and expected total cost of collisions.
 1 Introduction
-In the wakeup problem, devices are classified as either active (awake) or inactive (asleep).
-Each active device has a packet to send on a shared channel, while inactive participants are limited to listening.
-The first successful transmission acts as a “wakeup call” where all sleeping participants are woken up, and the problem is solved [2].
-Traditionally, prior work on this problem has focused on the time until the first success, which is typically referred to as the latency.
+In the wakeup problem, devices are either active (awake) or inactive (asleep).
+Each active device has a packet to send on a shared channel, while inactive devices are limited to listening.
+The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem [2].
+Prior work on this problem has focused on the time until the first success, which is typically referred to as the latency.
 Minimizing this metric is challenging, since in any unit of time, the shared channel can only accommodate a single sender; if two or more devices send simultaneously, the result is a collision where none of the corresponding packets are successfully sent.
-Moreover, sending on the channel is performed in a distributed fashion; that is, a priori there is no central scheduler or coordinator.
+Sending on the channel is performed in a distributed fashion; that is, a priori there is no central scheduler or coordinator.
 Under the standard cost model for wakeup, each collision wastes a single slot.
-However, in many settings, a collision can cost us much more than a single slot.
-For instance, in mobile systems, a failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range [26].
-For WiFi networks, a collision is known to consume time proportional to the packet size [1].
+However, in many settings, a collision can consume much more than a single slot.
 Therefore, collisions may significantly impact the performance of modern systems.
-In this work, we address both latency and collision cost in solving the wakeup problem.
-As we will discuss later in Section 3.1, while existing algorithms and natural approaches yield strong latency guarantees, they often result in high collision costs.
-Therefore, optimizing for these two metrics requires a new approach.
+In this work, we address both latency and the cost of collisions required to solve the wakeup problem.
+While existing algorithms give strong latency guarantees, they result in a large number of collisions (see Section 3.1), and addressing these two metrics in tandem requires new ideas.
+Why Should We Care About Collisions?
+The delay incurred by a collision can vary across different systems.
+As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consumes time, say $\mathcal{C}$, proportional to the packet size [1].
+In this setting, wakeup incurs an unavoidable delay of $\mathcal{C}$, since a single success is required.
+Yet, reducing the delay caused by collisions is still important.
+If this delay is, say, $X$, then throughput is $\mathcal{C}/(\mathcal{C}+X)$.
+So, if $X=\Theta(\mathcal{C})$, then throughput is sub-optimal by a constant factor.
+However, if say $X=\Theta(\sqrt{\mathcal{C}})$, then throughput is $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, throughput approaches the optimal value of $1$ as $\mathcal{C}$ grows large.
+[^1] Empty slots will also degrade throughput, but in WiFi, these slots are negligible compared to the time of sending a packet or a collision.
+This illustrates how minimizing collisions improves performance.
+More generally, over long-range links, where the propagation delays are large, a collision means devices can wait a long round-trip time before realizing the transmission has failed.
+Also, in intermittently-connected mobile wireless networks [21], failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range.
+Finally, in environments where there is some fluctuation in the delay (e.g., arising from fluctuating channel conditions or periods of heavy network usage), $\mathcal{C}$ can be interpreted as an upper bound on this delay that holds under most operating conditions.
 1.1 Model and Notation
@@
 However, if multiple packets transmit simultaneously, they all fail—this is a collision—and they may try again later.
-There is no ability to distinguish between an empty slot and a collision; that is, there is no collision detection.
-There are no identifiers for the packets.
-Finally, packets must be sent in a distributed manner; that is, there is no scheduler or central authority.
-Packet Activation.
-We consider two settings.
-For the static setting, in the first slot, all $n$ packets are active on the multiple access channel and begin executing an algorithm.
-By implication, all packets agree on the index of the current slot; this feature is typically referred to as a global clock [2].
-In the dynamic setting, all $n$ packets become active over time.
-To capture the worst case, we consider these activation times to be scheduled by an adversary, and we often refer to packets being activated or injected by the adversary.
-Packets may not agree on the index of the current slot, and thus there is no global clock.
+There is no scheduler or central authority a priori.
+No Collision Detection.
+There is no mechanism to distinguish between an empty slot and a collision—that is, no collision detection (no-CD).
+In many practical systems, collision detection (CD) is unavailable or unreliable.
+For example, in WiFi networks, received signal strength may offer a weak form of CD, but it is error-prone.
+Consequently, many prior results address the challenging no-CD setting (see the survey [2] and related discussion in Section 2).
+We emphasize that CD and the cost of collisions are orthogonal concerns: collisions incur a performance penalty regardless of whether they are detectable.
+For example, in WiFi, even if devices cannot sense an ongoing collision, the channel remains unavailable for its duration.
 Metrics.
 We care about latency, which refers to the time between the first packet injection and the first success.
-Additionally, each collision incurs a known cost of $\mathcal{C}$ $\geq 4$.
+Additionally, each collision incurs a known delay of $\mathcal{C}$ $\geq 4$.
 [^2] The value $4$ is chosen to simplify the analysis.
-We also note that only $\mathcal{C}$ $\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [16, 13, 2]) for the standard model will be asymptotically optimal.
+We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [13, 10, 2]) for the standard model will be asymptotically optimal.
 To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown.
 Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
-Given any wakeup algorithm, the pertinent costs are: (i) the collision cost, which is the number of collisions multiplied by $\mathcal{C}$, and (ii) the latency.
-Although both metrics represent time-based costs and we aim to minimize their maximum, we report them separately for clarity and to facilitate our lower-bound results.
-[…]
-For a random variable $X$, we say that w.h.p.
-in $x$ the expected value $E[X]\leq v$ if, when tallying the events in the expected value calculation, the total probability of those events where $X>v$ is $O(1/{\texttt{poly}({x})})$, where the constant degree of the polynomial is can be made arbitrarily large.
-Finally, our results are parameterized by $\epsilon$, which is a small positive constant $\ll 1$ set by our algorithm prior to execution.
-For terms of the form $\log^{k+(k^{\prime}/\epsilon)}n$, where $k,k^{\prime}$ are positive constants at least $1$, we report this as $\log^{\Theta(1/\epsilon)}n$ for ease of presentation.
+Given any wakeup algorithm, the two pertinent costs are: (i) the total cost of collisions, which is the number of collisions incurred over the execution of the algorithm multiplied by $\mathcal{C}$, and (ii) the latency.
+We report on both of these metrics separately, with the aim of minimizing their maximum.
+Notation.
+An event holds with high probability (w.h.p.) in $x$ if it occurs with probability at least $1-O(1/{\texttt{poly}({x})})$, where we can make the constant degree of the polynomial arbitrarily large.
+We use $\log^{c}y$ to denote $(\log y)^{c}$.
 1.2 Our Results
-Our algorithm, Aim-High (AH), applies to the static and dynamic wakeup problems.
-[^2] We explain the name in Section 3.1 We first address the static setting, which is closely tied to the classic problem of leader election (see Section 2).
-Here, we define $\eta$ $\max\{n,\mathcal{C}\}$, which is used when bounding the error of our upper bounds derived in Section 4.1.
+The performance of our algorithm, Aim-High (AH), is parameterized by $\epsilon$, which is a small positive constant set by the user prior to execution.
 Theorem 1.1
-W.h.p. in $\eta$, AH solves the static wakeup problem. The following associated costs hold w.h.p. in $\eta$:
+AH solves the wakeup problem with the following costs:
   1. 1.
-    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+2\epsilon}\right)$.
+    If $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
   2. 2.
-    Else, the latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
-By way of comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1).
-In contrast, Theorem 1.1 implies that if $\mathcal{C}$ is “large”, AH offers performance that is significantly sublinear in $\mathcal{C}$.
-Otherwise, the latency and the expected collision cost are each polylogarithmic-in-$n$, which echoes prior wakeup results without collision detection (see Section 2).
-Many upper bounds in the static setting are achieved by “fair” algorithms.
+    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
+By comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1).
+In contrast, Theorem 1.1 implies that if $\mathcal{C}$ is “large”, AH offers performance that is sublinear in $\mathcal{C}$.
+Else, latency and expected total cost of collisions are polylogarithmic-in-$n$, which echoes prior wakeup results (see Section 2).
+Many upper bounds are achieved by “fair” algorithms.
 A fair algorithm is defined by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability used by every packet in the $i$-th slot.
-In Section 4.2, we give the following lower bound for this class of algorithms (to which AH belongs) having a w.h.p.
-bound on the expected latency.
+In Section 5, we establish the following trade-off between latency and expected total cost of collisions.
 Theorem 1.2
 Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$.
-Any fair algorithm that w.h.p.
-in $\eta$ solves the static wakeup problem with $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has w.h.p.
-an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-k\epsilon}/\log^{\beta}n)$.
-This theorem, along with noting that the bound in (2) of Theorem 1.1 is within a polylogarithmic-in-$n$ factor of the optimal latency and expected collision cost, yields the following result.
-Corollary 1
-When $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, AH is within $O(\mathcal{C}^{4\epsilon})$-factor of the optimal expected collision cost for fair algorithms, and within a polylogarithmic-in-$n$ factor of the optimal latency and expected collision cost otherwise.
-In Section 4.3, we note that while AH upper bounds the latency and expected collision cost with w.h.p.
-in $\eta$, if this guarantee fails, then these quantities may be unbounded.
-We describe revisions to AH that yield the same bounds for expected latency and expected collision cost that hold with probability $1$.
-Next, we analyze the performance of AH in the dynamic setting, and we give the following upper-bound results. Here, our w.h.p. guarantees are in terms of $\mathcal{C}$.
-Theorem 1.3
-W.h.p. in $\mathcal{C}$, AH solves the dynamic wakeup problem. The following associated costs hold w.h.p. in $\mathcal{C}$:
-  1. 1.
-    If  $\mathcal{C}=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$ for a small constant $\delta>0$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+\epsilon}\log n\right)$,
-  2. 2.
-    Else, the latency is $\tilde{O}(n(\log^{\Theta(1/\epsilon)}n))$ and the expected collision cost is $\tilde{O}(n(\log^{\Theta(1/\epsilon)}n))$.
-Again, if $\mathcal{C}$ is large, then we achieve performance that is sublinear in $\mathcal{C}$.
-We give a lower bound for a class of algorithms we call “batch-fair" (to which AH belongs) that is analogous to fair algorithms, but natural in the dynamic setting.
-Theorem 1.4
-When $\mathcal{C}=\Omega(n^{2-\delta})$ for a small constant $\delta>0$, AH is within an $O(\mathcal{C}^{2\epsilon}\log^{2}n)$-factor of the optimal expected collision cost for batch-fair algorithms, and within an $O(\log^{\Theta(1/\epsilon)}n)$-factor of the optimal latency otherwise.
+Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
 2 Related Work
-The wakeup problem, has been extensively studied for decades. Here, we summarize closely-related results.
-Static Wakeup.
-In the literature, static wakeup, also known as leader election, addresses the fundamental problem of symmetry breaking.
-Early work by Willard [25] gives an upper bound of $\lg\lg n+O(1)$, and a matching lower bound for fair algorithms with collision detection (CD).
+Wakeup, or leader election, addresses the fundamental problem of symmetry breaking.
+Early work by Willard [20] establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
 In contrast, without CD, a common theme is that the latency increases to logarithmic or polylogarithmic in $n$.
@@
 A complementary by Jurdziński and Stachowiak [14] demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
-Given access to $b>1$ multiple channels with CD, Fineman et al.
-[12] give an algorithm with $O((\log n/\log b)+(\log\log n)(\log\log\log n))$ latency whp for any number of active devices.
+With multiple channels and collision detection, Fineman et al.
+[9] give an efficient contention-resolution algorithm.
 In the SINR model [17], Fineman et al.
-[11] demonstrate $O(\log n+\log R)$ latency w.h.p., where $R$ is the ratio of longest to shortest link, implying $O(\log n)$ latency as typically $R=O({\texttt{poly}({n})})$; they also provide a general $\Omega(\log n)$ lower bound.
-Chang et al [7] focus on energy efficiency under different models of channel feedback.
-Gilbert et al.
-[14] analyze the impact of system-size predictions, deriving lower and upper bounds.
-Dynamic Wakeup.
-The static and dynamic wakeup problems differ significantly.
-Specifically, for dynamic wakeup without CD, Jurdzinski and Stachowiak [19] show an $\Omega(\log^{2}n)$ lower bound, which contrasts with the $\Omega(\log n)$ lower bound [21] and matching upper bound [3] in the static setting.
-Without CD, Jurdziński and Stachowiak [18] explore the implications of global and local clocks, known or unknown $n$, packet labels, both with and without CD.
-Most relevant to our work is an algorithm that succeeds with probabilty at least $1-\varepsilon$ and has $O(n/\log n)$ latency when there is no knowledge of $n$, no global clock, no CD, and no labels; the probability of failure for the corresponding algorithm is parameterized
-Under the SINR model without a global clock, assuming devices know a polynomial upper bound on $n$, Jurdzinski and Stachowiak [19] give a Monte Carlo algorithm with $O(\log^{2}n/\log\log n)$ latency w.h.p., and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency, compared to a deterministic $O(\log^{2}n)$ latency [17].
-Without CD and no global clock, Chlebus et al.
-[9] explore the impact of multiple $b>1$ channels, where they give a randomized algorithm with error $\varepsilon$ and latency $O(n^{1/b}\ln(1/\varepsilon))$.
-Finally, we highlight prior work that addresses the cost of collisions in the context of the closely-related problem of contention resolution (see [8, 2]), where the goal is to have all $n$ packets successfully sent .
-Here, Biswas et al.
-[5] solve the static setting of contention resolution with latency and expected collision cost $\tilde{O}(n\sqrt{\mathcal{C}})$; however, they do not address the dynamic setting.
-Furthermore, Biswas et al.
-[5] assume CD, which their algorithm relies on to adjust the sending probability; this additional channel feedback is unavailable in our model.
-Another related result is by Anderton et al.
-[1] who show that collisions significantly degrade the performance of popular backoff algorithms for contention resolution.
-3 Algorithm Specification
+[8] achieve logarithmic latency and establish matching lower bounds.
+Chang et al.
+[6] examine energy efficiency under varying feedback models, while Gilbert et al.
+[11] study the effect of system-size prediction, deriving corresponding bounds.
+Research on the cost of collisions is still in its early stages.
+Anderton et al.
+[1] show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs.
+Building on this, Biswas et al.
+[4] provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
 1
@@
 4 $w_{0}\leftarrow 2^{\mathcal{C}^{\epsilon}}$
+5 $j\leftarrow 1$
+6
+7
+8// Iterations
+9 while true do
+10
+11 // Halving Phase
 12 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
-6 // Halving Phase
 13 while $w_{\mbox{\tiny cur}}\geq 2$ do
-8 for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ell$ do
-9
-10   Send packet with probability $1/w_{cur}$
+14    for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ln(w_{\mbox{\tiny cur}})$ do
+15
+16
+17     -2ptSend packet with probability $1/w_{cur}$
 18    if success then
@@
 21       $w_{\mbox{\tiny cur}}\leftarrow w_{\mbox{\tiny cur}}/2$
-15
-16// Doubling Phase
-17 while True do
+22
+23
+24 // Doubling Phase
+25 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
+26 for $s=1$ to $2^{j}$ do
 27    $w_{\mbox{\tiny cur}}\leftarrow 2w_{\mbox{\tiny cur}}$
-19 for slot $i=1$ to $d\ell$ do
-20
-21   Send packet with probability $1/w_{cur}$
+28    for slot $i=1$ to $d\sqrt{\mathcal{C}}\,\ln(w_{\mbox{\tiny cur}})$ do
+29
+30
+31     -2ptSend packet with probability $1/w_{cur}$
 32    if success then
 33        Terminate
-24
+34
+35 $j\leftarrow j+1$
 Algorithm 1 Aim-High
-The pseudocode for AH is given above, and each packet executes this algorithm upon being activated.
-Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$; recall that $\epsilon$ $\ll 1$ is a small positive constant that influences our bounds on cost.
-For the static setting, $\ell$ $\ln(w_{\mbox{\tiny cur}})$, where $w_{\mbox{\tiny cur}}$ is the current window size. For the dynamic setting, $\ell=\ln(\mathcal{C})$.
-Initially, the current window size $w_{\mbox{\tiny cur}}$ is set to $w_{0}$, and this is updated throughout the execution of two phases, as we describe below and illustrated in Figure 1.
+3 Algorithm Specification
+The pseudocode for AH is given above, and each packet executes this algorithm upon being activated. Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$.
+[…]
 […]
@@
 Doubling phase.
-At the beginning of this phase, the window size doubles such that now $w_{\mbox{\tiny cur}}=4$ (Line 1).
-Then, each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of $d\ell$ slots (Line 1 and 1); we refer to these slots as a small sample.
-If a success occurs, the execution terminates; otherwise, this process continues.
+At the beginning of this phase, $w_{\mbox{\tiny cur}}$ starts at $2w_{0}$.
+Each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of a sample (Line 1 and 1).
+If a success occurs, the execution terminates; otherwise, the window doubles and another sample is executed.
+This process continues until $2^{j}$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
 3.1 Context and Overview
-In this section, we start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
+We start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
 3.1.1 A Baseline via Backoff.
-In the static setting, a natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$.
+A natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$.
 At $i=\lg(n)$, the probability of a success in a slot becomes constant.
-This approach is a common ingredient in algorithm design; indeed, AH makes use of it.
-Another example of its use in solving wakeup is by Jurdziński and Stachowiak [18], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours, but without a notion of collision cost.
-However, in this prior result, each active packet starts with a constant sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
+An example of its use in solving wakeup is given by Jurdziński and Stachowiak [14], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
 So, we can observe that while backoff may be an important part of any solution, a key obstacle is how to set the initial sending probability suitably low.
-We do not know $n$ a priori, and so it seems that no matter what initial sending probability we choose—even if it depends on the known value of $\mathcal{C}$—the adversary may choose a sufficiently large value of $n$ such that collisions are unavoidable.
+Since $n$ is unknown a priori, it seems that no matter what initial sending probability we choose—even if it depends on the known value of $\mathcal{C}$—the adversary may choose a sufficiently large value of $n$ such that collisions are unavoidable.
 What can we do?
-Figure 1: Illustration of the phases for AH.
-The top portion depicts the halving phase, where the initial window size is $w_{0}=2^{\mathcal{C}^{\epsilon}}$.
-The bottom portion depicts the doubling phase, starting from a window of size $4$, where the degree to which packets contend for channel access decreases as the window grows, leading to a likely success.
 3.1.2 Design Choices and Overview.
 We now discuss how the design of AH allows us to do better than prior results in a model with collision cost.
-Setting an Initial Window Size. The initial sending probability needs to be set carefully. Too high and we incur many (potentially expensive) collisions as we have seen above.
-Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as a good window.
-Informally, once packets reach this good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$.
+Setting an Initial Window Size. The initial sending probability needs to be set carefully. Too high and we incur many collisions as we discussed above.
+Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as the size of a good window.
+Informally, once packets reach a good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$.
 Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success.
 At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots.
-This observation dictates the size of our large samples (recall Section 3).
-In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above informal analysis yields good expected collision cost (and, as we will show, good latency).
-However, since the adversary sets $\mathcal{C}$ with knowledge of $n$, this is by no means guaranteed; the adversary can always make collisions overwhelmingly likely.
-However, in this case, our consolation prize is that collisions should be cheap.
-To elaborate, we choose to "aim high" with the initial window size; that is, setting $w_{0}=2^{\mathcal{C}^{\epsilon}}$.
-This choice sets up our analysis of the following two cases for the static setting:
+This observation dictates the size of our samples, which are $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ and appear in lines 1 and 1 in the halving phase, and lines 1 and 1 in the doubling phase of our pseudocode.
+In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above reasoning yields good expected collision cost (and, as we will show, good latency).
+However, since the adversary sets $\mathcal{C}$ with knowledge of $n$, our hopes may be dashed; the adversary can always make collisions overwhelmingly likely.
+However, in this case, our consolation prize is that collisions are cheap, and the degree of cheapness corresponds to $w_{0}$.
+We choose to "aim high" with the initial window size; that is, setting $w_{0}=2^{\mathcal{C}^{\epsilon}}$.
+This choice sets up our analysis of the following two cases:
   1. 1.
-    Go Left For Good Window.
-If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good window exists at or to the left of the initial window; that is, the good-window size is less than or equal to the initial-window size.
-Here, the probability of a collision is “not too high”.
+    If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good-window size is less than or equal to the initial-window size.
+Here, the probability of a collision is not too high.
 Thus, so long as we adjust the sending probability upwards in a conservative manner, we can reduce expected collision cost; this is the purpose of the halving phase.
-Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the time until the first success will be too large.
+Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the latency will be too large.
   2. 2.
-    Go Right For Good Window Conversely, if $n\sqrt{\mathcal{C}}>w_{0}$, then the good window exists to the right of the initial window.
+    Else $n\sqrt{\mathcal{C}}>w_{0}$, and the good-window size is larger than the initial-window size.
 Here, the collision probability may be “too high”.
-However, a key realization here is that the collision cost is cheap, since we will argue that $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$ implies that $\mathcal{C}=O(\log^{1/\epsilon}n)$.
+However, a key realization is that $\mathcal{C}$ is small; we will argue that $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$ implies that $\mathcal{C}=O(\log^{1/\epsilon}n)$.
 Thus, we can tolerate these collisions as we increase the window size in the doubling phase.
-While the above cases address the particulars of our approach to the static setting, analogous intuition holds for the dynamic version (see Section 5.1 for discussion).
-Samples.
-As discussed above, we "aim high" with the initial window and execute the the halving phase; here, we are hoping to succeed via samples of size $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$.
-However, if the adversary chooses to set $\mathcal{C}$ such that $n\sqrt{\mathcal{C}}>w_{0}$, good contention will not be discovered by the halving phase, and our ambitious setting of $w_{0}$ does not yield a success in the halving phase.
-In this case, however, collisions are now provably cheap, and we no longer care about finding a good window; we simply want a quick success, even if we rack up collisions in the process.
-Thus, the doubling phase—which is classic backoff—uses smaller samples of size $d\log(w_{\mbox{\tiny cur}})$, and once a window of size $\Theta(n)$ is reached, a success is likely.
-Finally, we note that in the static setting, the start and end points of each sample are the same for all packets.
-However, for the dynamic setting, this may not hold given that the adversary controls when packets are activated; that is, the samples of packets can be unsynchronized, and this poses an additional challenge.
+Why Iterations?
+Although AH is likely to terminate when a good window is reached, there is some small probability of failure.
+Suppose we are in Case 1, but this failure occurs in the halving phase.
+Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 3).
+However, if the halving phase is never revisited, the packets may accrue intolerable latency by executing the doubling phase without success (as the sending probability drops off exponentially fast).
+Conversely, suppose we are in Case 2 and this failure occurs.
+Then, if we never reset $w_{\mbox{\tiny cur}}$ to (roughly) $w_{0}$, the latency costs may again be high.
+Therefore, iterations facilitate good expected latency (see Lemma 6).
 Parameter $\epsilon$.
 Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis.
-Yet, it is worth pointing out how it interacts with our metrics of interest: latency and expected collision cost.
+Yet, it is worth pointing out how it interacts with the latency and the expected total cost of collisions.
 The former is improved by making $\epsilon$ small, since this shortens the halving phase.
@@
 However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$.
-Thus, degree of the polynomial increases with the reciprocal of $\epsilon$, which means our expected collision cost increases.
-4 The Static Setting
-In this section, we analyze the performance of AH for the static wakeup problem. Throughout our analysis, due to space constraints, some proofs have been moved to the appendix.
-4.1 Upper Bound
-Throughout, we assume that $n$ is sufficiently large. We make use of the following well-known facts.
+Thus, degree of the polynomial increases with the reciprocal of $\epsilon$.
+4 Upper Bound
+In this section, we analyze the performance of AH.
+Define $\eta$ $\max\{n,\mathcal{C}\}$.
+A good window was loosely discussed in Section 3.1.2; we define it as any window with size in $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$.
+Throughout, we assume that $n$ is sufficiently large.
+We make use of the following well-known facts.
 Fact 1
@@
 […]
-4.1.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
-In this case, we show that w.h.p.
-AH achieves a success during the halving phase.
-We refer to a window with size in the range $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ as good.
-We first bound the latency required to reach a good window, and then analyze the expected collision cost.
+The next lemma argues that success is likely in a good window.
 Lemma 1
-Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, AH terminates at or before a good window is reached.
-Lemma 2
-Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, the latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
+W.h.p. in $\eta$, a success occurs if AH executes in a good window.
 Proof
-Under AH, each packet starts with an initial window size of $w_{0}=2^{\mathcal{C}^{\epsilon}}$.
-By assumption, the good window size is at most $w_{0}$, and so a good window will be reached during the halving phase.
-The halving phase executes, and by Lemma 1, w.h.p.
-in $\eta$ we are guaranteed to succeed at or before a good window.
-The number of windows that AH executes over before reaching a good window is at most $\lg(w_{0})=\mathcal{C}^{\epsilon}$, and the sample size for each such window is $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}}))\leq d\sqrt{\mathcal{C}}C^{\epsilon}$.
-Therefore, the latency is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
+Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3). The probability of a success in any fixed slot of this good window is at least:
+[…]
+$\geq\left(\frac{1}{3e\sqrt{\mathcal{C}}}\right)\mbox{~~~since $\mathcal{C}\geq 4$.}$
+Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$.
+The expected number of successes in the sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots is at least:
+$E[X]\geq\left(\frac{d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})}{3e\sqrt{\mathcal{C}}}\right)$ $=\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$
+where $d^{\prime}$ is a constant that can be made arbitrarily large depending on $d$.
+We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
+$Pr\left(X\leq(1-\delta)\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$ $\leq Pr\left(X\leq\frac{d\ln(w_{\mbox{\tiny cur}})}{6e}\right)$
+$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
+$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
+$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
+$\leq\eta^{-d^{\prime\prime}}$
+where the third line follows from $w_{\mbox{\tiny cur}}=\Theta(\ln n\sqrt{\mathcal{C}})$, and $d^{\prime}$ and $d^{\prime\prime}$ can be arbitrarily large constants for $d$ is sufficiently large.
+4.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
+In this case, the good window size is less than or equal to $w_{0}$, so we argue that AH is likely to obtain a success during the halving phase.
+We make use of the following known result that upper bounds the probability of a collision.
 Lemma 2
@@
 Lemma 3
-Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, the expected collision cost for AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
+Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected total cost of collisions for AH is $O(\mathcal{C}^{(1/2)+\epsilon})$.
 Proof
-In the halving phase, by Lemma 1, w.h.p.
-in $\eta$ a success occurs at or before a good window, which we are guaranteed by assumption to reach since $n\sqrt{\mathcal{C}}\leq w_{0}$.
-Thus, we consider any window of size $w\geq n\sqrt{\mathcal{C}}$ (and no more than $w_{0}$).
-We upper bound the probability of a collision using Lemma 3, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$).
-Consider all the slots from the samples, and index these slots from $i\geq 1$.
-Let $X_{i}$ be an independent random variable where $X_{i}=1$ when the $i$-th slot contains a collision; $X_{i}=0$ otherwise.
-Then, by Lemma 3, we have:
-$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}.$
-Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ be the size of the first good window that is reached during halving phase.
-To get an upper bound on the expected collision cost, we first sum up the probabilities of a collision in any of the windows from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to $w^{*}\geq n\sqrt{\mathcal{C}}$, which requires at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$ halvings.
-Then, we multiply by the total number of slots in all the samples until a good window is reached, and then multiply by the cost of a collision.
-To start, the sum of the probabilities is as follows:
-$\sum_{j=0}^{C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2}\hskip-8.0pt\frac{2\left(\frac{n}{w_{0}/2^{j}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{j}}\right)\right)}$ $=\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\lg\mathcal{C})}{2}}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{j}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{j}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
-$\leq\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\ln\mathcal{C})}{2}}\hskip 0.0pt\frac{2^{2j}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$
-$\leq\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\ln\mathcal{C})}{2}}\hskip 0.0pt2^{2j}$
-$=\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$
-$<\frac{8}{\mathcal{C}}.$
-We note that the second line by assumption of $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$. The third line holds for any $\mathcal{C}\geq 4$.
-Recall that the sample size is $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=O(\mathcal{C}^{(1/2)+\epsilon})$.
-Over at most $\lg(w_{0})=O(C^{\epsilon})$ samples, the total number of slots in all the samples is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
-Thus, letting $X=\sum_{i}X_{i}$ and using our upper bound on the probability above, the expected number of collisions is :
-$E\left[X\right]$ $\leq E\left[\sum_{i=1}^{O(\mathcal{C}^{(1/2)+2\epsilon})}\hskip-8.0ptX_{i}\right]=O\left(\frac{\mathcal{C}^{(1/2)+2\epsilon}}{\mathcal{C}}\right)$
-Thus, multiplying this by $\mathcal{C}$, we obtain that the expected collision cost is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
-4.1.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
-In this case, we show that w.h.p. in $\eta$ AH succeeds during the doubling phase. To begin, we show that $\mathcal{C}$ is small in $n$.
+[…]
+[…]
+$G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ (1)
+Upper Bounding $B_{j}$.
+Clearly, $B_{j}$ is upper-bounded by the cost of having a collision in every single slot of iterations $1$ through $j$.
+In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\ln(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contributes $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots.
+Additionally, in iteration $k$, the corresponding doubling phase executes $2^{k}$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, by the sum of a geometric series, over iterations $1$ through $j$, the doubling phases contributing $O(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j})$ slots.
+By the above:
+$B_{j}$ $=O\left(j\,d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j}\right)\cdot\mathcal{C}$
+$=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\log(w_{0})2^{j}\right)\cdot\mathcal{C}$
+$=O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
+since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(2^{j}w_{0})=O(j\mathcal{C}^{\epsilon})$.
+By Lemma 1, the $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ where we can make the constant $c>0$ as large as we wish. Thus,
+$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
+$=O(1)$
+where the last line follows for sufficiently large $c$.
+Upper Bounding $G$.
+To upper bound $G$, for each slot $i$ in the first halving phase of the first iteration, let $X_{i}$ be an independent random variable where $X_{i}=1$ if the $i$-th slot contains a collision; otherwise, $X_{i}=0$.
+Recall that the windows of this first halving phase are repeatedly halved.
+For a slot in a window of size $w=w_{0}/2^{k}$, for a positive integer $k\geq 0$, by Lemma 2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$), we have:
+$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}$
+$=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}$
+The expected cost due to collisions that occur in this particular window is thus at most:
+$Pr(X_{i}=1)\cdot d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\cdot\mathcal{C}$ $\leq\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}d\mathcal{C}^{(3/2)+\epsilon}.$
+Note that the number of windows executed starting from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to size $n\sqrt{\mathcal{C}}$, is at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$; for ease of presentation, let $x=C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$.
+By the above, we can derive:
+$G$ $\leq\sum_{k=0}^{x}\hskip 0.0pt\frac{2\left(\frac{n}{w_{0}/2^{k}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{k}}\right)\right)}\left(d\mathcal{C}^{(3/2)+\epsilon}\right)$
+$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\sum_{k=0}^{x}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
+$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$ (2)
+$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$ (3)
+$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$ (4)
+$<\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{8}{\mathcal{C}}$
+$=O\left(\mathcal{C}^{(1/2)+\epsilon}\right).$
+In the above, line 2 holds since $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$, since we are in Case 1.
+Line 3 holds for any $\mathcal{C}\geq 4$, and line 4 holds by the sum of the geometric series.
+Thus, Equation 1 is bounded as claimed.
+The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears.
+Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
+Lemma 4
+Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
+4.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
+In this case, the good window size exceeds $w_{0}$, and so we argue that AH likely succeeds during the doubling phase. A key insight is that, for Case 2, $\mathcal{C}$ is small in terms of $n$.
 Lemma 5
@@
 as desired, since $\epsilon$ is a constant.
-Next, we show argue that a success is likely once we reach a window of size $\Theta(n)$.
+The next lemma establishes bounds on expected latency and total collision cost, and the intuition is as follows.
+In Case 2, collisions are “cheap” by Lemma 5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$.
+Thus, the cost of collisions incurred until the good window is reached is small; it takes only $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=O(\log n)$ doublings of the window size to reach a good window in the doubling phase.
+Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by $j=O(\log\log n)$.
+After this point the probability of success is high, and so subsequent iterations contribute little to the expected total cost of collisions.
 Lemma 6
-Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, in the doubling phase, AH achieves a success at or before a window of size $n$ is reached.
+Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+4}n)\right)$.
+Proof
+The number of slots incurred for each individual halving phase is the sample size times the number of halvings, so upper bounded by:
+$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(2^{j}w_{0})\log(w_{0})\right)$
+$=O\left(\sqrt{\mathcal{C}}j\mathcal{C}^{2\epsilon}\right)$
+$=O\left(j\log^{1/(2\epsilon)+2}n\right)$
+where the last line follows by applying Lemma 5. Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
+The number of slots in the doubling phase of the $j$-th iteration is:
+$2^{j}d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O(2^{j}\sqrt{\mathcal{C}}\log(2^{j}w_{0}))$
+$=O(2^{j}\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon})$
+$=O(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n))$
+slots, where the last line follow by Lemma 5. By the above, summing up the number of slots executed over iterations 1 to $j$ is $O(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n))$.
+Pessimistically, let us assume no success ever occurs during the halving phase.
+Focusing on the doubling phase, a good window is reached within at most $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=\log_{2}n+O(\log\log n)$ doublings of the window size.
+Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log\log n$, where $c^{\prime}>0$ is a sufficiently large constant.
+By the above, the total number of slots up until a good window is reached is:
+$O\left(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log n)\,(\log\log n)^{2}(\log^{1/(2\epsilon)+2}n)\right)$
+$=O\left(\log^{1/(2\epsilon)+4}n\right)$
+After this point the probability of success is high, by Lemma 1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
+$O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
+$=O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{n}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-10.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
+$=O\left(\log^{1/(2\epsilon)+4}n\right)$
+which completes the argument.
+The following lemma about the expected total cost of collisions is proved similarly to Lemma 6.
+The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
 Lemma 7
-Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, the latency required by AH is $O(\log^{4+(1/(2\epsilon))}n)$.
+Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+4}n\right)$.
+4.3 Performance of AH
+We can now prove our upper bounds on the latency and expected collision cost for AH.
+Theorem 1.1. AH solves the wakeup problem with the following costs:
+  1. 1.
+    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
+  2. 2.
+    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
 Proof
-Under AH, all packets start with an initial window size of $w_{0}$, and they iteratively halve the window size until a window size of $2$ (in the worst case) is reached, or until a success, whichever occurs first.
-Pessimistically, assume no success occurs during the halving phase.
-By Lemma 2, the latency incurred for the halving phase is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
-After that, the doubling phase begins.
-W.h.p.
-in $\eta$, by Lemma 6, a success at or before a window of size $n$; this requires $O(\log n)$ doublings and, for each window size, we have a sample of $O(\log(w_{\mbox{\tiny cur}}))=O(\log n)$ Therefore, the latency for the doubling phase is $O(\log^{2}n)$.
-It follows that the sum of the latency from the halving phase and the doubling phase is:
-$O(\mathcal{C}^{(1/2)+2\epsilon}\log^{2}n)$ $=O(\log^{4+(1/(2\epsilon))}n)$
-since $\mathcal{C}=O(\log^{1/\epsilon}n)$ by Lemma 5.
-Lemma 8
-Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, the collision cost is $O(\log^{4+(3/(2\epsilon))}n)$.
-Proof
-By assumption $n\sqrt{\mathcal{C}}>w_{0}$, so we can invoke Lemma 7, which states that w.h.p.
-in $\eta$ the latency is $O(\log^{4+(1/(2\epsilon))}n)$.
-Pessimistically, every slot may contain a collision, and by Lemma 5, $\mathcal{C}=O(\log^{1/\epsilon}n)$.
-Thus, an upper bound on the total collision cost is:
-$O(\mathcal{C}\,\log^{4+(1/(2\epsilon))}n))$ $=O(\log^{4+(3/(2\epsilon))}n)$
-which yields the claim.
-4.1.3 Total Cost of AH
-We can now prove our upper bounds on the latency and expected collision cost for AH in the static setting.
-Theorem 1.1. W.h.p. AH solves the static wakeup problem with the following costs:
-  1. 1.
-    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+2\epsilon}\right)$.
-  2. 2.
-    Else, the latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ latency and the expected collision cost is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
-Proof
-There are two cases to consider.
-In the first case, $n\sqrt{\mathcal{C}}\leq w_{0}$.
-By Lemmas 2 and 4, w.h.p.
-the latency and expected collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
+The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are there are two cases to consider.
+In Case 1, $n\sqrt{\mathcal{C}}\leq w_{0}$.
+By Lemmas 3 and 4, the expected latency and the expected total collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
 In the second case, $n\sqrt{\mathcal{C}}>w_{0}$.
-By Lemmas, 7 and 8, w.h.p.
-when $n\sqrt{\mathcal{C}}>w_{0}$, AH has $O(\log^{4+(1/(2\epsilon))}n)$ latency and the collision cost is $O(\log^{4+(3/(2\epsilon))}n)$, yielding (2).
-Finally, we note that the relationship between $n$ and $\mathcal{C}$— a relationship established by the adversary (recall Section 1.1)—dictates which upper bound applies.
-Since (1) dominates (2) when $C=\Omega(\log^{\Theta(1/\epsilon)}n)$, the result follows.
-4.2 Lower Bound
+By Lemmas, 6 and 7, the expected latency and expected total cost of collisions are $O(\log^{4+\frac{1}{2\epsilon}}n)$ and $O(\log^{4+\frac{3}{2\epsilon})}n)$, respectively, yielding (2).
+5 Lower Bound
 […]
 Lemma 8
-Any fair algorithm for the static wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
-[…]
-Lemma 10 (Biswas et al. [6], Lemma 24)
-Fix any slot $t$ in any fair algorithm for the static wakeup problem. If $\texttt{Con}(t)>2$, then the probability of a collision is at least $1/10$.
-By Lemma 10, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$.
-Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$.
-Under these circumstances, we provide a lemma that lower bounds the probability of a collision as a function of contention.
-We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
-Lemma 11
-Fix any slot $t$ where there are at least $2$ packets and $\texttt{Con}(t)\leq 2$ under the execution of any fair algorithm for the static wakeup problem.
-The probability of a collision in $t$ is $\Omega(\texttt{Con}^{2}(t))$.
-Proof
-For a fair algorithm, in any slot $t$, $\texttt{Con}(t)=np$.
-For any fixed slot, the probability of a collision involving exactly two packets (since there are at least $2$ packets by assumption) is a lower bound on the probability of a collision (that may involve more than $2$ packets), and this is $\binom{n}{2}p^{2}(1-p)^{n-2}$, which is at least:
-$\geq\binom{n}{2}p^{2}(1-p)^{n}\geq\left(\frac{n^{2}}{2}\right)\left(\frac{\texttt{Con}(t)}{n}\right)^{2}e^{-2\cdot\texttt{Con}(t)}\geq\left(\frac{\texttt{Con}^{2}(t)}{2e^{2\cdot\texttt{Con}(t)}}\right)$
-$=\Omega(\texttt{Con}^{2}(t))$
-where the last line follows from $\texttt{Con}(t)\leq 2$.
-We can derive a relationship between latency and expected collision cost for fair algorithms.
-Theorem 1.2.
-Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$.
-Any fair algorithm that w.h.p.
-in $\eta$ solves the static wakeup problem with $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has w.h.p.
-an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-k\epsilon}/\log^{\beta}n)$.
-Proof
-Let $\mathcal{A}$ be an algorithm that w.h.p.
-in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is any constant.
-By Lemma 9, $\mathcal{A}$ has total contention:
-$\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$ (2)
-As we discussed, an implication of Lemma 10, we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot.
-Thus, we can use Lemma 11 (noting that $n\geq 2$), which implies that the expected collision cost in a single slot $t$ is:
-[…]
-for some constant $k>0$. So, the total expected collision cost is:
-$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\mathcal{C}$
-By Jensen’s inequality for convex functions, we have:
-$\frac{\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)}{L}\geq\left(\frac{\sum_{t=1}^{L}k\,\texttt{Con}(t)}{L}\right)^{2}$
-which implies:
-$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\geq\frac{\left(\sum_{t=1}^{L}k\,\texttt{Con}(t)\right)^{2}}{L}$ (3)
-So, by Equation 3, the expected cost is at least:
-[…]
-and since w.h.p., $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$, we have:
-$\frac{\mathcal{C}\,k^{2}\,2^{2}}{L}$ $=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
-which yields the claim.
-Recalling Theorem 1.1, if the adversary selects $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then Theorem 1.2 implies that AH is within $O(\mathcal{C}^{4\epsilon})$-factor of the optimum expected collision cost for fair algorithms.
-For smaller values of $\mathcal{C}$, we are clearly within a polylogarithmic-in-n factor of the optimal latency and expected collision cost.
-These observations yield our Corollary 1.
-4.3 Improved Bounds on Cost of Aim-High
-AH achieves its bounds on latency and expected collision cost with high probability in $\eta$.
-However, note that if this guarantee fails, these quantities may be unbounded.
-This can happen in two ways.
-First, say that $n\sqrt{\mathcal{C}}\ll w_{0}$ and AH does not succeed in a good window.
-Then, the packets will complete the halving phase and execute the doubling phase, where the probability of a success can be vanishingly small.
-Second, when $n\sqrt{\mathcal{C}}>w_{0}$ and AH is unlucky and does not succeed within a window of size $\Theta(n)$, the packets may execute the doubling phase in perpetuity.
-Intuitively, any fix requires periodically giving the packets another chance to succeed, for both cases: $n\sqrt{\mathcal{C}}\leq w_{0}$ and $n\sqrt{\mathcal{C}}>w_{0}$.
-Consequently, we modify AH in the following way.
-Modifying AH.
-The new algorithm runs in iterations indexed by $j\geq 0$.
-For iteration $j=0$, the packets execute an entire halving phase.
-If there is no success, the packets then execute a truncated doubling phase; they run the original doubling phase for $2^{j}=1$ small samples (or until a success) starting at window size $4$; after this, iteration $j=0$ has ended.
-In the next iteration $j=1$, the packets execute another entire halving phase, and then execute another truncated doubling phase for $2^{j}=2$ small samples, again starting at window size $4$.
-And the process continues until the first success.
-Intuition.
-If $n\sqrt{\mathcal{C}}\leq w_{0}$, then we now have a new chance of success in the halving phase of each iteration.
-Since success occurs w.h.p.
-in $\eta$, failing in the good window is such a low-probability event that the expected cost of the next $2^{j}$ small samples in the following truncated doubling phase is small.
-Conversely, if $n\sqrt{\mathcal{C}}>w_{0}$, then the halving is useless, and we are searching for a window of size $\Theta(n)$ in the following truncated doubling phase of the iteration.
-Since $\mathcal{C}$ is $O(\log^{1/\epsilon}n)$ in this case, we can afford the extra latency and collision cost from the interleaved halving phases.
-Moreover, consider when an iteration is reached where the truncated doubling phase has packets increasing their window size to $\Theta(n)$; that is, once the truncated doubling phase runs for $\lg n$ samples (or, in other words, $j=\lg\lg n$).
-Then, the contribution to the expected cost from the next truncated doubling phase is made small by the low probability of failing in the current truncated doubling phase.
-The proofs for both statements below are in the appendix.
-Lemma 12
-Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency and expected collision cost are $O(\mathcal{C}^{(1/2)+2\epsilon})$.
-Lemma 13
-Assume $n\sqrt{\mathcal{C}}>w_{0}$. The expected latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
-5 The Dynamic Setting
-[…]
-5.1 Technical Overview
-Compared to the static setting, there are some significant differences in our analysis for the dynamic setting.
-For starters, there is no good window, given that multiple batches may be active, each with their respective windows.
-Instead, we have the analogous notion of good contention, which is contention in the range $3/\sqrt{\mathcal{C}},27/\sqrt{\mathcal{C}}]$.
-In Lemma 15, we show that a success is likely if we have $\Omega(\sqrt{\mathcal{C}})$ slots with good contention.
-At time slot $t$, contention is said to high when $\texttt{Con}(t)>27/\sqrt{\mathcal{C}}$, and low when $\texttt{Con}(t)<3/\sqrt{\mathcal{C}}$.
-Unlike the static setting, we cannot divide our analysis into two cases (i.e., $n\sqrt{\mathcal{C}}\leq w_{0}$ and $n\sqrt{\mathcal{C}}>w_{0}$) quite as cleanly.
-For instance, in the dynamic setting, suppose that $n\sqrt{\mathcal{C}}>w_{0}$ and the adversary initially injects a small number of packets $n^{\prime}\leq w_{0}/\sqrt{\mathcal{C}}$ which execute the halving phase.
-  • $\bullet$
-    If no more packets are activated, these $n^{\prime}$ packets may reach good contention and yield a success.
-This is despite $n\sqrt{\mathcal{C}}>w_{0}$, which would not have occurred in the static setting.
-  • $\bullet$
-    Else, suppose the adversary injects the remaining packets just prior to these $n^{\prime}$ packets reaching good contention.
-All of sudden, the contention jumps spikes (preventing a success) in a manner that was not possible in the static setting.
-[…]
-Given that the sending probabilities of packets in separate batches will differ, a technical challenge arises when trying to bound the probability of a collision.
-We tackle this via an exchange argument in the proof of Lemma 16, and then in Lemma 17 we leverage this to bound the collision cost as a function of contention.
-A similar challenge is addressed in trying to show that good contention can be reached during the halving phase for $t^{*}$ sufficiently large.
-Again, the dynamic aspect of the problem prevents a straightforward analysis.
-Thus, we argue via Lemma 18 that, if there had been high contention in the halving phase, we would been able to “rewind” backwards in the execution and show that there was a period of good contention leading to a success and termination; in other words, we can derive a contradiction.
-Error.
-In the static case, a success occurs when all $n$ packets have a window large enough such that using $\ell=\ln(w_{\mbox{\tiny cur}})$ in the sample size yields guarantees that are w.h.p.
-in $\eta$.
-However, in the dynamic setting, this is not the case; a success may occur when the currently-active packets (in an assortment of batch sizes) have small respective windows.
-Thus, in this setting, AH succeeds with probability at least $1-O(\mathcal{C}^{-d^{\prime}})$, where $d^{\prime}>0$ is a constant that can be made arbitrarily large depending only on sufficiently large $d$.
-5.2 Upper Bound
-Throughout, we index slots starting at $1$, which corresponds to the slot where the first packet is injected; that is, we ignore all possible empty slots prior to the first packet.
-Since the minimum window size is $2$ under the execution of AH, the largest sending probability is $1/2$, and so we make use of the following prior result for both cases $t\leq t^{*}$ and $t>t^{*}$.
-Lemma 14
-(Bender et al [4]) Assume for any fixed slot $t$ that the sending probability of each active packet is at most $1/2$.
-Then, the probability of a success in slot $t$ is at least $\texttt{Con}(t)/e^{2\texttt{Con}(t)}$.
-5.2.1 Case 1: $t\leq t^{*}$.
-To show w.h.p. in $\mathcal{C}$ a success occurs, we will need the following analog to Lemma 1.
-Lemma 15
-After at least $d\sqrt{\mathcal{C}}\ln\mathcal{C}$ slots with good contention, w.h.p. in $\mathcal{C}$ a packet has succeeded.
-Our analysis of collision cost for this case depends on the following. For $n$ probability values $p_{1},p_{2},...,p_{n}$, consider the following:
-$S=\sum_{i\neq j\neq\cdots\neq k}\overbrace{p_{i}p_{j}\dots p_{k}}^{\alpha}$
-which is the sum of all products of $\alpha$ probability values for $1\leq\alpha\leq n$. The next lemma addresses the maximum value of this sum.
-Lemma 16
-There is a setting of the $n$ probability values to be equal that maximizes $S$.
-Proof
-Consider any setting of the probability values that maximizes $S$.
-If this setting already has all the probability values set equal, then we are done.
-Else, there exist two indices $i$ and $j$ such that $p_{i}<p_{j}$.
-We now present an exchange argument.
-Consider a small shift of $\delta=\frac{p_{j}-p_{i}}{2}$ from $p_{j}$ to $p_{i}$, keeping all other probability values the same.
-That is, let ${p_{i}}\leftarrow p_{i}+\delta$ and ${p_{j}}\leftarrow p_{j}-\delta$.
-This redistribution affects any terms in $S$ that involve $p_{i}$ or $p_{j}$.
-Consequently, we analyze its impact by considering two cases: terms that involve both $p_{i}$ and $p_{j}$, and terms that involve only one of them.
-At first, we consider the change in $S$ for the first case. Each summand in $S$ involving both $p_{i}$ and $p_{j}$ changes as follows:
-$(p_{i}+\delta)(p_{j}-\delta)\cdots p_{k}=(p_{i}p_{j}-\delta p_{i}+\delta p_{j}-\delta^{2})\cdots p_{k}.$
-Focusing on the first factor on the right, we have:
- $\mbox{\hskip 8.0pt}(p_{i}p_{j}-\delta p_{i}+\delta p_{j}-\delta^{2})$
-$=p_{i}p_{j}-\left(\frac{p_{j}-p_{i}}{2}\right)p_{i}+\left(\frac{p_{j}-p_{i}}{2}\right)p_{j}-\left(\frac{p_{j}-p_{i}}{2}\right)^{2}$
-$=\frac{1}{4}\left(4p_{i}p_{j}-2p_{i}p_{j}+2p^{2}_{i}+2p^{2}_{j}-2p_{i}p_{j}-p^{2}_{j}-p^{2}_{i}+2p_{i}p_{j}\right)$
-$=\frac{1}{4}\left(p^{2}_{j}+p^{2}_{i}+2p_{i}p_{j}\right)$
-where the last line is strictly positive.
-Now, for the second case, note that for every summand with $p_{i}$ and not $p_{j}$, there is a “matching” term with $p_{j}$ and not $p_{i}$, and with all other probability terms in the summand the same.
-Thus, we analyze the increase due to $(p_{i}+\delta)p_{l}\cdots p_{k}$ and the decrease from the matching term $(p_{j}-\delta)p_{l}\cdots p_{k}$, and show that the net result is at least 0.
-Formally, is the following true?
-$\left((p_{i}+\delta)p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)\hskip-1.0pt-\hskip-1.0pt\left(p_{i}p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)\hskip-1.0pt\geq\hskip-1.0pt\left(p_{j}p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)\hskip-1.0pt-\hskip-1.0pt\left((p_{j}-\delta)p_{l}\cdot\hskip-2.0pt\cdot\hskip-2.0pt\cdot p_{k}\right)$
-Observe that by canceling out a factor of $p_{l}\cdots p_{k}$, the inequality is true.
-Thus, by the above analysis, the net change is positive, which contradicts our assumption that $S$ was already maximized.
-Packets in different batches have different sending probabilities, which complicates our analysis. To this end, the next lemma captures the probability of a collision as a function of contention.
-Lemma 17
-For any slot with low or good contention, the probability of a collision is at most $O(1/\mathcal{C})$.
-Proof
-We establish this argument by comparing collisions involving different numbers of packets.
-Collisions may occur between pairs, triples, quadruples, and so on.
-The probability of an $\alpha$-way collision—a collision involving any $\alpha\geq 2$ packets—is at most
-$\leq\sum_{i\neq j\neq\dots\neq k}\overbrace{p_{i}p_{j}\cdots p_{k}}^{\alpha}$
-$\leq\sum_{1}^{\binom{n}{\alpha}}(p^{*})^{\alpha}$
-$=\binom{n}{\alpha}\left(\frac{27}{n\sqrt{\mathcal{C}}}\right)^{\alpha}$
-$=O\left(\frac{1}{\mathcal{C}^{\alpha/2}}\right)$ (4)
-The second line follows from Lemma 16, where $p^{*}$ denotes the common sending probability of all packets.
-The third line follows from the definition of the upper bound of good $\texttt{Con}(t)=27/\sqrt{\mathcal{C}}$ (which, of course, is an upper bound on low contention).
-In any fixed slot, we compute the probability of collisions involving exactly two packets by setting $\alpha=2$ in Equation 4, which yields a bound of $O(1/\mathcal{C})$.
-Similarly, the total probability of collision involving three or more packets ($\alpha\geq 3$) is:
-$O\left(\sum_{\alpha=0}^{n}\left(\frac{1}{\sqrt{\mathcal{C}}}\right)^{\alpha}\right)$ $=O\left(\sum_{\alpha=3}^{n}\left(\frac{1}{\sqrt{\mathcal{C}}}\right)^{\alpha}-1-\frac{1}{\sqrt{\mathcal{C}}}-\frac{1}{\mathcal{C}}\right)$
-$=O\left(\frac{1}{\mathcal{C}^{3/2}}\right)$
-Therefore, the total probability of collisions at most $\Theta(1/\mathcal{C})+O(1/(\mathcal{C}^{3/2}))=O(1/\mathcal{C})$.
-Recall that for $t\leq t^{*}$, $n(t)\leq w_{0}/\sqrt{\mathcal{C}}$ and, here, we wish to argue that a success will occur while packets are executing their respective halving phases.
-Note that the contention can be good or low in the initial window; pessimistically, consider the case where it is low.
-Moreover, at some point in the halving process, as the windows shrink in size, contention will be high.
-We need to argue that, in between these points of low and high contention, there is a period of good contention.
-Our next lemma provides a tool for showing this is true; this is done by considering a point in time in the halving phase where contention can be high (or good) and “rewinding” backwards in the execution, while tracking contention.
-Let s $d\sqrt{\mathcal{C}}\ell=d\sqrt{\mathcal{C}}\ln\mathcal{C}$ denote the large sample size used during the halving phase of in AH (recall Section 3).
-Lemma 18
-Suppose at time $t\leq t^{*}$, all active packets are executing the halving phase, and $\texttt{Con}(t)\geq 3/\sqrt{\mathcal{C}}$. Then:
-$C(t-s)\in\left[\frac{\texttt{Con}(t)}{3},\frac{\texttt{Con}(t)}{2}\right].$
-Proof
-Rewinding by $s$ slots from the current time $t$, each batch that exists at time $t$ either (1) still exists back at time tst-s, but with windows twice as large, or (2) does not exist because it has not yet been injected back at time tst-s.
-In the case of (1), since all batches exist back at time tst-s with their respective windows doubled, then $C(t-s)\leq\texttt{Con}(t)/2$.
-Else, in the case of (2), some packets that are present at time $t$ are not yet injected at time tst-s.
-To measure the change in contention, we ask: how much contention is present at time $t$, but absent back at time tst-s due to these packets not yet active?
-Note that such packets can only have been in the system for at most $s$ slots (otherwise, they would still be present at time tst-s); therefore, each of these packets has a window size of $w_{0}$.
-Since by assumption $n(t)\leq w_{0}/\sqrt{\mathcal{C}}$, this means that the contention contributed by all such packets at time $t$ is at most $n(t)/w_{0}\leq 1/\sqrt{\mathcal{C}}$.
-Thus, we can get a lower bound on the contention at time tst-s by first subtracting out this contribution, and then accounting for the doubling of the windows for those packets that do exist back at time tst-s.
-That is, $C(t-s)\geq\frac{\texttt{Con}(t)-(1/\sqrt{\mathcal{C}})}{2}$.
-To simplif this , we note that $\frac{\texttt{Con}(t)-(1/\sqrt{\mathcal{C}})}{2}$ $\geq\texttt{Con}(t)/3$, since by assumption $\texttt{Con}(t)\geq 3/\sqrt{\mathcal{C}}$.
-[…]
-Lemma 19
-Suppose that $t^{*}>k\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C}$ for a sufficienty large constant $k>0$.
-Then, w.h.p.
-in $\mathcal{C}$, AH yields a success in $\mathcal{I}$ with latency and expected collision cost $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
-5.2.2 Case 2: $t>t^{*}$.
-We start with an analog to Lemma 5.
-Lemma 20
-Assume $t>t^{*}$. Then, $\mathcal{C}=O(\log^{1/\epsilon}n)$.
-Proof
-By definition of $t^{*}$, if $t>t^{*}$, then $n(t)>w_{0}/\sqrt{\mathcal{C}}$.
-Solving for $\mathcal{C}$ is done as in the proof of Lemma 5 yields $\mathcal{C}=O(\log^{1/\epsilon}n(t))$, and the result follows from noting that $n(t)\leq n$ for any slot $t$.
-We also need an analog to Lemma 6 (used in the static setting), which shows that $d\ell$ slots with constant contention (rather than a window of size $n$) yields a success with high probability in $\ell$.
-For this, we define the interval of adequate contention to be $[1/16,1]$.
-Lemma 21
-Over $d\ln\mathcal{C}$ slots of adequate contention, w.h.p. in $\mathcal{C}$, a success occurs.
-For a sufficiently large constant $\kappa>0$, define a gap to be a contiguous set of at least $\gamma$ = $\kappa(\log^{1+(1/(2\epsilon))}n)\log\log n$ slots, during which the adversary makes no packet injections.
-We note that the value of $\kappa$ is lower bounded in the proof of Lemma 22 below.
-Lemma 22
-Consider any time slot $t>t^{*}$ that marks the start of a gap. Then, w.h.p. in $\mathcal{C}$, a success occurs during the gap.
-Lemma 23
-W.h.p.
-in $\mathcal{C}$, a success occurs within $O(n(\log^{1+(1/(2\epsilon))}n)$ $\log\log n)$ slots of $t^{*}$ and with an associated $O(n(\log^{1+(3/(2\epsilon))}n)$ $\log\log n)$ collision cost over these slots.
-Proof
-Given a gap starting in any slot $t>t^{*}$, by Lemma 22 w.h.p.
-in $\mathcal{C}$, there is a success in this gap.
-Pessimistically, assume no success occurs before the first gap (otherwise, our latency and collision cost is smaller).
-Starting from slot $t^{*}+1$, what is the maximum number of slots that the adversary can force to elapse before the first gap must occur?
-To avoid a gap from occurring, a minimum of one packet must be activated at least every $\gamma-1$ slots.
-Given that there are at most $n$ packets ever activated, this implies that the number of slots prior to the first gap and a consequent success is at most $n\gamma$ $O(n(\log^{1+(1/(2\epsilon))}n)\log\log n)$ slots.
-The collision cost is upper bounded by considering a collision in every slot, which is:
-$O(\mathcal{C}n(\log^{1+(1/(2\epsilon))}n)\log\log n)$ $=O(n(\log^{1+(3/(2\epsilon))}n)\log\log n)$
-by Lemma 20, which completes the argument.
-5.2.3 Total Cost of AH
-We can now prove bounds on latency and the expected collision cost of AH for the dynamic setting.
-Theorem 1.3. W.h.p. in $\mathcal{C}$, AH solves the dynamic wakeup problem with the following costs:
-  1. 1.
-    If  $\mathcal{C}=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$ for a small constant $\delta>0$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+\epsilon}\log n\right)$,
-  2. 2.
-    Else, the latency is $O(n(\log^{1+(1/(2\epsilon))}n)\log\log n)$ and the expected collision cost is $O(n(\log^{1+(3/(2\epsilon))}n)\log\log n)$.
-Proof
-If $t^{*}>k\mathcal{C}^{(1/2)+\epsilon}\log\mathcal{C}$, for a sufficiently large constant $k>0$, Lemma 19 implies that w.h.p.
-in $\mathcal{C}$, a success occurs with latency and expected collision cost $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
-Else, by Lemma 23, w.h.p.
-in $\mathcal{C}$, a success occurs within $O(n$ $(\log^{1+(1/(2\epsilon))}n)$ $\log\log n)$ slots of $t^{*}$ and with an associated $O(n$ $(\log^{1+(3/(2\epsilon))}n)$ $\log\log n)$ collision cost over these slots.
-What about the additional latency and collision cost up to slot $t^{*}$?
-By Lemmas 19 and 20, each of these is:
-$O(\mathcal{C}^{(1/2)+\epsilon}\log\mathcal{C})=O((\log^{1+(1/(2\epsilon))}n)\log\log n)$
-which is asymptotically smaller, and so the claimed bounds follow.
-Finally, given as input a (large) value of $n$ to the algorithm, the adversary may choose which upper bound applies for any given execution by setting $\mathcal{C}$ and $t^{*}$.
-For sufficiently large $\mathcal{C}$, the bounds in (1) dominate the bounds in (2).
-Specifically, this occurs for:
-$\mathcal{C}$ $=\Omega(n^{1/(0.5+\epsilon)}\log^{\Theta(1/\epsilon))}n)=\Omega(n^{2-\delta}\log^{\Theta(1/\epsilon)}n)$
-where the last line follows for some small constant $\delta>0$.
-We note that the costs stated in (1) of Theorem 1.3 differ from those in the corresponding (1) of Theorem 1.1.
-This occurs because of the different values of $\ell$ (recall Section 3).
-In the dynamic case, $\ell=\ln\mathcal{C}$, which is used to get our w.h.p.
-in $\mathcal{C}$ guarantees, and is the source of the $O(\log n)$-factor.
-In contrast, for the static case $\ell=\ln(w_{\mbox{\tiny cur}})$, which is used to get our w.h.p.
-in $\eta$ guarantees, and is the source of the additional $O(\mathcal{C}^{\epsilon})$-factor.
-5.3 Lower Bound
-In the dynamic setting, Jurdziński and Stachowiak [18] highlight the class of algorithms where all packets exhibit identical behavior, governed by a sequence $p_{1}$, $p_{2}$, $\dots$, where $p_{i}$ denotes the sending probability in the $i$-th time slot after a packet is activated; we refer to this class of algorithms as batch-fair.
-Observe that, when restricted to the case where all packets start in slot $1$ (i.e., the static setting), a batch-fair algorithm is fair. Thus, we can apply Theorem 1.2 here to obtain the following.
-Lemma 24
-Any batch-fair algorithm that w.h.p. solves dynamic wakeup with $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$ latency must incur an expected collision cost of $\Omega(\mathcal{C}^{(1/2)-\epsilon}/\log n)$.
-Note that AH is a batch-fair algorithm for the dynamic case.
-By Theorem 1.3, if the adversary selects $C$ sufficiently large—that is, $\mathcal{C}=\Omega(n^{2-\delta})$—then Lemma 24 implies that AH is within an $O(\mathcal{C}^{2\epsilon}\log^{2}n)$-factor of the optimal expected collision cost.
-For smaller values of $\mathcal{C}$, however, Lemma 24 is not useful.
-In this case, we consider a prior result by Jurdziński and Stachowiak [18], who show a lower bound on latency of $\Omega(n/\log n)$ for batch-fair algorithms, which applies to our setting.
-Therefore, for smaller values of $\mathcal{C}$, AH is within an $O(\log^{\Theta(1/\epsilon)}n)$-factor of the optimal latency.
-These observations yield Theorem 1.4.
-6 Future Work
-For the dynamic setting, an interesting open problem is whether we can achieve finite expected latency and collision cost bounds that hold with probability $1$.
-We would also like to prove general lower bounds for the static and dynamic settings.
-Finally, it may be worthwhile to explore resilience to jamming under this model.
-Appendix
-Here, we provide arguments that could not be included in the main paper due to space constraints.
-Proofs from Section 4.1
-Lemma 1. Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, AH terminates at or before a good window is reached.
-Proof
-If a success occurs before a good window, then the claim is true.
-Else, we are at a good window with size $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3), which occurs within $O(\log w_{0})$ halvings.
-Here, the probability of a success in any fixed slot is at least:
-[…]
-$\geq\left(\frac{1}{3e\sqrt{\mathcal{C}}}\right)\mbox{~~~since $\mathcal{C}\geq 4$.}$
-Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$.
-The expected number of successes in the sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots is at least:
-$E[X]\geq\left(\frac{d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})}{3e\sqrt{\mathcal{C}}}\right)$ $=\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$
-We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
-$Pr\left(X\leq(1-\delta)\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$ $\leq Pr\left(X\leq\frac{d\ln(w_{\mbox{\tiny cur}})}{6e}\right)$
-$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
-$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
-$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
-$\leq\eta^{-d^{\prime\prime}}$
-where the third line follows from $w_{\mbox{\tiny cur}}=\Theta(\ln n\sqrt{\mathcal{C}})$, and $d^{\prime}$ and $d^{\prime\prime}$ can be arbitrarily large constants for $d$ is sufficiently large.
-Lemma 6. Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, in the doubling phase, AH achieves a success at or before a window of size $n$ is reached.
-Proof
-Suppose we are in a window of size $w=n$. The probability of a success in any fixed slot is at least:
-[…]
-$\geq 1/e^{2}.$
-Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$. The expected number of successes in the sample over $d\ln(w_{\mbox{\tiny cur}})$ is at least:
-$E[X]\geq\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{e^{2}}\right).$
-We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
-$Pr\left(X\leq(1-(1/2))\frac{d\ln(w_{\mbox{\tiny cur}})}{e^{2}}\right)$ $\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)e^{2}}}$
-$=e^{-\frac{d\ln(n)}{10e^{2}}}$
-$\leq n^{-d^{\prime}}$
-$\leq\eta^{-d^{\prime\prime}}$
-The second follows from $w_{\mbox{\tiny cur}}=\Theta(n)$, and where $d^{\prime\prime}$ and $d^{\prime}$ can be arbitrarily large constant depending only on $d$ being sufficiently large.
-The last line follows from Lemma 5, which implies that $n\gg\mathcal{C}$.
-Proofs from Section 4.2
-Lemma 9. Any fair algorithm for the static wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
+Any fair algorithm for the wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
 Proof
@@
 Thus, there is a constant probability of all packets failing, which contradicts the w.h.p. guarantee of $\mathcal{A}$.
-Proofs from Section 4.3
-Lemma 12. Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency and expected collision cost are $O(\mathcal{C}^{(1/2)+2\epsilon})$.
+[…]
+Lemma 9(Biswas et al. [5], Lemma 24)
+Fix any slot $t$ in any fair algorithm for the wakeup problem. If $\texttt{Con}(t)>2$, then the probability of a collision is at least $1/10$.
+By Lemma 9, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$.
+Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$.
+Under these circumstances, we provide a lemma that lower bounds the probability of a collision as a function of contention.
+We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
+Lemma 10
+Fix any slot $t$ where there are at least $2$ packets and $\texttt{Con}(t)\leq 2$ under the execution of any fair algorithm for the wakeup problem.
+The probability of a collision in $t$ is $\Omega(\texttt{Con}^{2}(t))$.
 Proof
-In the halving phase of iteration $j$, the latency and collision cost is $O(\mathcal{C}^{(1/2)+2\epsilon})$ by Lemmas 2 and 4.
-In the truncated doubling phase of iteration $j$, the latency is
-$O(2^{j}\ell)$ $=O(2^{j}\log(w_{\mbox{\tiny cur}}))$
-$=O(2^{j}\log(2^{j}))$
-$=O(j2^{j})$
-and the corresponding collision cost is at most $O(j2^{j}\mathcal{C})$, since in the worst case there is a collision in every slot; note that this dominates the latency.
-By Lemma 1, for any fixed halving phase, the probability of failing in a good window is $O(1/{\texttt{poly}({\eta})})$, where we can make the degree of the polynomial an arbitrarily large constant depending on sufficiently large $d$.
-Thus, each of the expected latency and collision cost is at most: the bound from Case (1) of Theorem 1.1 for the w.h.p.
-in $\eta$ event of succeeding in iteration $j=0$, plus the increasingly low-probability event of failing up until the $j$-th iteration for $j\geq 1$ times the associated cost.
-That is:
-$=O\left(\mathcal{C}^{(1/2)+2\epsilon}\right)+\sum_{j=1}^{\infty}O\left(\frac{1}{{\texttt{poly}({\eta})}}\right)^{j}O\left(j2^{j}\mathcal{C}\right)$
-$=O\left(\mathcal{C}^{(1/2)+2\epsilon}\right)$
-which completes the argument.
-Lemma 13. Assume $n\sqrt{\mathcal{C}}>w_{0}$. The expected latency and expected collision cost are $O(\log^{\Theta(1/\epsilon)}n)$.
+For a fair algorithm, in any slot $t$, $\texttt{Con}(t)=np$.
+For any fixed slot, the probability of a collision involving exactly two packets (since there are at least $2$ packets by assumption) is a lower bound on the probability of a collision (that may involve more than $2$ packets), and this is $\binom{n}{2}p^{2}(1-p)^{n-2}$, which is at least:
+$\geq\binom{n}{2}p^{2}(1-p)^{n-2}$
+$\geq\binom{n}{2}p^{2}(1-p)^{n}\geq\left(\frac{n^{2}}{2}\right)\left(\frac{\texttt{Con}(t)}{n}\right)^{2}e^{-2\cdot\texttt{Con}(t)}\geq\left(\frac{\texttt{Con}^{2}(t)}{2e^{2\cdot\texttt{Con}(t)}}\right)$
+$=\Omega(\texttt{Con}^{2}(t))$
+where the last line follows from $\texttt{Con}(t)\leq 2$.
+We can derive a relationship between latency and expected collision cost for fair algorithms.
+Theorem 1.2.
+Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$.
+Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
 Proof
-The latency and collision cost for the halving phase of iteration $j$ are each $O(\mathcal{C}^{(1/2)+2\epsilon})$ by Lemmas 2 and 4, and this becomes $O(\log^{O(1/\epsilon)}n)$ by Lemma 5.
-For the truncated doubling phase of iteration $j$, the latency is
-$O(2^{j}\ell)$ $=O(2^{j}\log(w_{\mbox{\tiny cur}}))$
-$=O(2^{j}\log(2^{j}))$
-$=O(j2^{j})$
-The corresponding collision cost is at most $O(j2^{j}\mathcal{C})=O(j2^{j}\log^{1/\epsilon}n)$, since in the worst case there is a collision in every slot, and we can invoke Lemma 5; note that this dominates the latency.
-A window of size $n$ is reached by at most the $\lg\lg(n)$-th iteration; $2^{j}$ samples are executed, each sample doubles the window, and we need only $\lg(n)$ doublings.
-Thus, the latency and collision cost up to a window of size $n$ is $O((\log^{1/\epsilon}n){\texttt{poly}({\log n})})=O(\log^{\Theta(1/\epsilon)}n)$.
-By Lemma 6, in a window of size $n$, the probability of failure is $O(1/{\texttt{poly}({\eta})})$. Using Case (2) of Theorem 1.1, the expected cost is this:
-$=O(\log^{\Theta\left(\frac{1}{\epsilon}\right)}n)+\hskip-5.0pt\sum_{j=\lg\lg n}^{\infty}O\left(\frac{1}{{\texttt{poly}({\eta})}}\right)^{j-\lg\lg(n)+1}O\left(j2^{j}\log^{\frac{1}{\epsilon}}n\right)$
-$=O(\log^{\Theta(1/\epsilon)}n)$
-which completes the argument.
-Proofs From Section 5.2
-Lemma 15. After at least $d\sqrt{\mathcal{C}}\ln\mathcal{C}$ slots with good contention, w.h.p. in $\mathcal{C}$ a packet has succeeded.
-Proof
-By the definition of good contention, $3/\sqrt{\mathcal{C}}\leq\texttt{Con}(t)\leq 27/\sqrt{\mathcal{C}}$. Thus, by Lemma 14, we have that the probability of a success in slot $t$ is:
-$\frac{\texttt{Con}(t)}{e^{2\texttt{Con}(t)}}$ $=\frac{b}{\sqrt{\mathcal{C}}}.$
-for some constant $b>0$. The probability of failure over $d\sqrt{\mathcal{C}}\ln{n}$ slots is:
-$\left(1-\frac{b}{\sqrt{\mathcal{C}}}\right)^{d\sqrt{\mathcal{C}}\ln{\mathcal{C}}}$ $\leq e^{-d^{\prime}\ln{\mathcal{C}}}$
-$\leq\frac{1}{\mathcal{C}^{d^{\prime}}}$
-where the first line follows from Fact 1 (a), and the constant $d^{\prime}$ is can be made arbitrarily large depending on sufficiently large $d$.
-Lemma 22. Consider any time slot $t>t^{*}$ that marks the start of a gap. Then, w.h.p. in $\mathcal{C}$, a success occurs during the gap.
-Proof
-Let $B$ be the most-recent batch injected prior to the gap starting at $t$.
-At some later slot $t^{\prime}$, the packets in $B$ will have completed the halving phase; this requires a number of slots at most:
- $\mbox{\hskip 10.0pt}s\lg(w_{0})$
-$=(d\sqrt{\mathcal{C}}\ln\mathcal{C})\lg(w_{0})$
-$=O((\log^{1/(2\epsilon)}n)(\log\log n)\mathcal{C}^{\epsilon})$
-$=O((\log^{1/(2\epsilon)}n)(\log\log n)(\log n))$
-$=O((\log^{1+(1/(2\epsilon))}n)\log\log n)$
-using Lemma 20.
-Since $B$ was the last batch injected prior to the gap starting, we know at slot $t^{\prime}$ that all active packets are executing the doubling phase.
-We also know that contention is high at slot $t^{\prime}$, since any packet in $B$ has a window of size $4$ (at the start of the doubling phase) and, thus, contributes at least $1/2$ to contention.
-During the doubling phase for any fixed batch, over each small sample of $\sigma=d\ln\mathcal{C}$ slots, the corresponding sending probability for those packets is halved.
-Due to different injection times, these small samples may be unsynchronized (recall Section 3.1.2).
-However, in every $s^{\prime}=2\sigma$ contiguous slots, all batches are guaranteed to halve their sending probability at least once and at most twice.
-Thus, starting at slot $t$, after every $s^{\prime}$ slots, the contention from packets in the system is reduced by a constant factor between $2$ and $4$.
-Given that adequate contention spans a factor of $16$ (recall, $[1/16,1]$), the contention will enter this range (i.e., it cannot be skipped), and this occurs within at most:
-$s^{\prime}O(\lg(\texttt{Con}(t)))$ $\leq O(s^{\prime}\lg(n))$
-$=O(\log n\log\log n)$
-slots, where the second line follows given that contention is always upper bounded by $n$ and Lemma 20.
-Additionally, again, given that the range of adequate contention spans a factor of $16$, at least one full small sample will execute with good contention; that is, contention will remain good for at least $d\ln\mathcal{C}$ slots.
-Therefore, by Lemma 21, w.h.p.
-in $\mathcal{C}$ there is a success.
-The total latency measured from slot $t$ is at most:
-$=O((\log^{1+(1/(2\epsilon))}n)\log\log n)+O(\log n\log\log n)$
-$\leq k(\log^{1+(1/(2\epsilon))}n)\log\log n$
-where $k$ is a sufficiently large positive constant. Thus, for $\kappa$ (in the definition of a gap) chosen such that $\kappa\geq k$, the claim holds.
-Lemma 19.
-Suppose that $t^{*}>k\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C}$ for a sufficienty large constant $k>0$.
-Then, w.h.p.
-in $\mathcal{C}$, AH yields a success in $\mathcal{I}$ with latency and expected collision cost $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
-Proof
-Since $n(t)\leq w_{0}/\sqrt{\mathcal{C}}$ for $t\leq t^{*}$, the contention in slot $1$ is $n(1)/w_{0}\leq 1/\sqrt{\mathcal{C}}$.
-Thus, at the start of the execution, contention can only be low or good (but not high).
-Next, we argue that within $\mathcal{I}$ the contention cannot change from low to high without good contention slots in between.
-Over any two consecutive slots in $\mathcal{I}$, the contention from active packets may double if all the corresponding windows halve.
-Also, again, the largest possible increase in contention due to newly-injected packets is $n(t)/w_{0}$ $\leq 1/\sqrt{\mathcal{C}}$, since $t\leq t^{*}$.
-Thus, given that good contention spans a factor of $9$ (recall $3/\sqrt{\mathcal{C}},27/\sqrt{\mathcal{C}}]$), contention cannot transition from low to high without at least some period of good contention in between.
-Do we have high contention in $\mathcal{I}$?
-Consider the first batch, $B$, injected in slot $1$.
-Barring a success, the packets in $B$ will execute their halving phase until a window of size $2$ is reached within $t^{\prime\prime}=O(s\log w_{0})=O(C^{(1/2)+\epsilon}\log\mathcal{C})$ slots, at which point they contribute high contention.
-By the above, this implies a period of good contention in $\mathcal{I}$.
-How many slots of good contention are in $\mathcal{I}$?
-There are two possibilities, but we will show that only the first can occur.
-In Possibility (1), at least $s=d\sqrt{\mathcal{C}}\ln\mathcal{C}$ slots of good contention occur before any slots with high contention.
-In Possibility (2), less than $s$ slots of good contention occur before any slots with high contention.
-For Possibility (1), Lemma 15 guarantees w.h.p.
-in $\mathcal{C}$ a success by the end of those $s$ slots of good contention and, as derived above, this occurs by slot $t^{\prime\prime}$.
-Thus, the latency is $O(\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C})$ $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$ given that $C={\texttt{poly}({n})}$ (recall Section 1.1).
-By Lemma 17, since contention up to this point is low or good, the probability of a collision up until this termination is $O(1/C)$, yielding $O((1/\mathcal{C})\mathcal{C}^{(1/2)+\epsilon}\ln\mathcal{C})$ collisions in expectation, and thus an expected collision cost of $O(\mathcal{C}^{(1/2)+\epsilon}\log\mathcal{C})$ $O(\mathcal{C}^{(1/2)+\epsilon}\log n)$.
-For Possibility (2), let $t^{\prime}\in\mathcal{I}$ be the first slot with high contention; this occurs before packets in $B$ finish their halving phase (since, again, $B$ contributes high contention when window size $2$ is reached).
-Since $B$ is the first batch injected, at $t^{\prime}$ all active packets are executing a halving phase.
-Thus, we can use Lemma 18.
-Starting at $t^{\prime}$, while contention remains at least $3/\sqrt{\mathcal{C}}$, for every $s$ slots we rewind into the past, the contention decreases by at least a factor of $2$ and at most a factor of $3$.
-We repeatedly invoke Lemma 18 in order to rewind to a point in time where contention is in $3/\sqrt{\mathcal{C}},9/\sqrt{\mathcal{C}}]$, which we claim occurs in $\mathcal{I}$.
-To see this, note that:
-$\texttt{Con}(t^{\prime})$ $\leq n(t^{\prime})/2\leq w_{0}/(2\sqrt{\mathcal{C}})$
-since $t^{\prime}\in\mathcal{I}$.
-Thus, dividing the maximum possible contention by $2$ at most $\mathcal{C}^{\epsilon}$ times lands us in $3/\sqrt{\mathcal{C}},9/\sqrt{\mathcal{C}}]$.
-That is, we must rewind by at most $\mathcal{C}^{\epsilon}$ samples – at most $dC^{(1/2)+\epsilon}\ln\mathcal{C}$ slots – which is accommodated by $t^{*}>kC^{(1/2)+\epsilon}\ln\mathcal{C}$, for $k$ sufficiently large.
-Finally, even if batches have unsynchronized samples (recall Section 3.1.2), we are guaranteed at least $s$ slots with contention in the range $3/\sqrt{\mathcal{C}},27/\sqrt{\mathcal{C}}]$.
-Thus, at least $s$ slots of good contention did occur within $\mathcal{I}$ prior to high contention, which contradicts Possibility (2).
-Lemma 21. Over $d\ln\mathcal{C}$ slots of adequate contention, w.h.p. in $\mathcal{C}$, a success occurs.
-Proof
-Using Lemma 14, the probability of success in any slot $t$ with adequate contention is $\texttt{Con}(t)/e^{2\texttt{Con}(t)}\geq 1/(16e^{1/8})>1/20$.
-The probability of no success over all $d\ln\mathcal{C}$ slots is at most:
-$(1-1/20)^{d\ln\mathcal{C}}$ $\leq e^{d\mathcal{C}/20}$
-$=\mathcal{C}^{-d^{\prime}}$
-where the first line follows by Fact 1(a), and where the constant $d^{\prime}>0$ can be made arbitrarily large for a sufficiently large $d$.
+Let $\mathcal{A}$ be an algorithm that w.h.p.
+in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is any constant.
+By Lemma 8, $\mathcal{A}$ has total contention:
+$\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$ (6)
+As we discussed, an implication of Lemma 9, we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot.
+Thus, we can use Lemma 10 (noting that $n\geq 2$), which implies that the expected cost of collisions in a single slot $t$ is:
+[…]
+for some constant $k>0$. So, the expected total cost of collisions is:
+$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\mathcal{C}$
+By Jensen’s inequality for convex functions, we have:
+$\frac{\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)}{L}\geq\left(\frac{\sum_{t=1}^{L}k\,\texttt{Con}(t)}{L}\right)^{2}$
+which implies:
+$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\geq\frac{\left(\sum_{t=1}^{L}k\,\texttt{Con}(t)\right)^{2}}{L}$ (7)
+So, by Equation 7, the expected cost is at least:
+[…]
+and since $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$, we have:
+$\frac{\mathcal{C}\,k^{2}\,2^{2}}{L}$ $=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
+which yields the claim.
+As evidence that our result is not far off from the asymptotic optimal, we note that any fair algorithms with latency $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ must have an expected total cost of collisions that is $\Omega(\mathcal{C}^{\frac{1}{2}-2\epsilon})$, which is an $O(\mathcal{C}^{3\epsilon})$-factor smaller than that achieved by AH.
+6 Conclusion and Future Work
+The theoretical exploration of collision costs remains in its early stages. Our work here provides steps towards addressing fundamental problems that connect with the wireless domain.
+For future work, we aim to prove more general lower bounds. Additionally, we wish to derive results for the dynamic setting, where packets arrive over time rather than all being initially present.

### v3

@@
 A Gentle Wakeup Call Symmetry Breaking with Less Collision Cost
-Abstract
+Abstract.
 The wakeup problem addresses the fundamental challenge of symmetry breaking.
@@
 The goal is to achieve a single successful transmission efficiently.
-Prior work on wakeup primarily analyzes latency—the number of slots until the first success.
+Prior work primarily analyzes latency—the number of slots until the first success.
 However, in many modern systems, each collision incurs a nontrivial delay, $\mathcal{C}$, which prior analyses neglect.
@@
 Finally, for a well-studied class of algorithms, we establish a trade-off between latency and expected total cost of collisions.
-1 Introduction
+1. Introduction
 In the wakeup problem, devices are either active (awake) or inactive (asleep).
 Each active device has a packet to send on a shared channel, while inactive devices are limited to listening.
-The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem [2].
+The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem (Banicescu et al., 2024).
 Prior work on this problem has focused on the time until the first success, which is typically referred to as the latency.
@@
 The delay incurred by a collision can vary across different systems.
-As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consumes time, say $\mathcal{C}$, proportional to the packet size [1].
+As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consume time, say some large value $\mathcal{C}$, proportional to the packet size (Anderton et al., 2021).
 In this setting, wakeup incurs an unavoidable delay of $\mathcal{C}$, since a single success is required.
 Yet, reducing the delay caused by collisions is still important.
-If this delay is, say, $X$, then throughput is $\mathcal{C}/(\mathcal{C}+X)$.
-So, if $X=\Theta(\mathcal{C})$, then throughput is sub-optimal by a constant factor.
-However, if say $X=\Theta(\sqrt{\mathcal{C}})$, then throughput is $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, throughput approaches the optimal value of $1$ as $\mathcal{C}$ grows large.
-[^1] Empty slots will also degrade throughput, but in WiFi, these slots are negligible compared to the time of sending a packet or a collision.
-This illustrates how minimizing collisions improves performance.
+If this delay is, say, $X$, then $\mathcal{C}/(\mathcal{C}+X)$ provides a measure of how much time is wasted due to collisions.
+So, if $X=\Theta(\mathcal{C})$, then this ratio is suboptimal by a constant factor.
+However, if $X=\Theta(\sqrt{\mathcal{C}})$, then $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, the ratio approaches the optimal value of $1$ as $\mathcal{C}$ grows large.
 More generally, over long-range links, where the propagation delays are large, a collision means devices can wait a long round-trip time before realizing the transmission has failed.
-Also, in intermittently-connected mobile wireless networks [21], failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range.
+Also, in intermittently-connected mobile wireless networks (Zhang, 2006), failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range.
 Finally, in environments where there is some fluctuation in the delay (e.g., arising from fluctuating channel conditions or periods of heavy network usage), $\mathcal{C}$ can be interpreted as an upper bound on this delay that holds under most operating conditions.
-1.1 Model and Notation
+1.1. Model and Notation
 […]
@@
 For example, in WiFi networks, received signal strength may offer a weak form of CD, but it is error-prone.
-Consequently, many prior results address the challenging no-CD setting (see the survey [2] and related discussion in Section 2).
+Consequently, many prior results address the challenging no-CD setting (see the survey (Banicescu et al., 2024) and related discussion in Section 2).
 We emphasize that CD and the cost of collisions are orthogonal concerns: collisions incur a performance penalty regardless of whether they are detectable.
@@
 [^1] The value $4$ is chosen to simplify the analysis.
-We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [13, 10, 2]) for the standard model will be asymptotically optimal.
-To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown.
+We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff (Greenberg and Leiserson, 1985; Geréb-Graus and Tsantilas, 1992; Banicescu et al., 2024)) for the standard model will be asymptotically optimal.
+To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree of the polynomial is unknown.
 Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
@@
 We use $\log^{c}y$ to denote $(\log y)^{c}$.
-1.2 Our Results
+1.2. Our Results
 The performance of our algorithm, Aim-High (AH), is parameterized by $\epsilon$, which is a small positive constant set by the user prior to execution.
-Theorem 1.1
+Theorem 1.1.
 AH solves the wakeup problem with the following costs:
-  1. 1.
+  1. (1)
     If $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
-  2. 2.
-    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
+  2. (2)
+    Else, the expected latency is $O(\log^{5+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{5+\frac{3}{2\epsilon}}n)$.
 By comparison, existing approaches to wakeup incur a collision cost that is $\Omega(\mathcal{C})$ (see Section 3.1).
@@
 In Section 5, we establish the following trade-off between latency and expected total cost of collisions.
-Theorem 1.2
+Theorem 1.2.
 Let $\alpha$ and $\beta$ be constants such that $1/(2\epsilon)\geq\alpha\geq 0$ and $\beta\geq 0$.
 Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
-2 Related Work
+2. Related Work
 Wakeup, or leader election, addresses the fundamental problem of symmetry breaking.
-Early work by Willard [20] establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
+Early work by Willard (Willard, 1986) establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
 In contrast, without CD, a common theme is that the latency increases to logarithmic or polylogarithmic in $n$.
-For starters, without collision detection and when $n$ is known, Kushilevitz and Mansour [16, 15] prove a lower bound of $\Omega(\log n)$ expected latency.
-Newport [18, 19], in work that unifies many prior lower-bound results, shows $\Omega(\log n)$ expected latency and $\Omega(\log^{2}n)$ latency w.h.p.
+For starters, without collision detection and when $n$ is known, Kushilevitz and Mansour (Kushilevitz and Mansour, 1998, 1993) prove a lower bound of $\Omega(\log n)$ expected latency.
+Newport (Newport, 2014a, b), in work that unifies many prior lower-bound results, shows $\Omega(\log n)$ expected latency and $\Omega(\log^{2}n)$ latency w.h.p.
 for general randomized algorithms.
 Complementing these lower bounds, Bar-Yehuda et al.
-[3] achieve a matching upper bound of $O(\log n)$ expected latency and w.h.p.
+(Bar-Yehuda et al., 1992) achieve a matching upper bound of $O(\log n)$ expected latency and w.h.p.
 $O(\log^{2}n)$ latency.
 Farach-Colton et al.
-[7] obtain $\Omega(\log n\log(1/\varepsilon))$ latency for algorithms without CD, succeeding with probability at least $1-\varepsilon$.
-A complementary by Jurdziński and Stachowiak [14] demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
+(Farach-Colton et al., 2006) obtain $\Omega(\log n\log(1/\varepsilon))$ latency for algorithms without CD, succeeding with probability at least $1-\varepsilon$.
+A complementary by Jurdziński and Stachowiak (Jurdziński and Stachowiak, 2002) demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
 With multiple channels and collision detection, Fineman et al.
-[9] give an efficient contention-resolution algorithm.
-In the SINR model [17], Fineman et al.
-[8] achieve logarithmic latency and establish matching lower bounds.
+(Fineman et al., 2016b) give an efficient contention-resolution algorithm.
+In the SINR model (Moscibroda and Wattenhofer, 2006), Fineman et al.
+(Fineman et al., 2016a) achieve logarithmic latency and establish matching lower bounds.
 Chang et al.
-[6] examine energy efficiency under varying feedback models, while Gilbert et al.
-[11] study the effect of system-size prediction, deriving corresponding bounds.
+(Chang et al., 2019) examine energy efficiency under varying feedback models, while Gilbert et al.
+(Gilbert et al., 2021) study the effect of system-size prediction.
 Research on the cost of collisions is still in its early stages.
 Anderton et al.
-[1] show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs.
+(Anderton et al., 2021) show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs.
 Building on this, Biswas et al.
-[4] provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
+(Biswas et al., 2024a) provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
 1
@@
 25 $w_{\mbox{\tiny cur}}\leftarrow w_{0}$
-26 for $s=1$ to $2^{j}$ do
+26 for $s=1$ to $j$ do
 27    $w_{\mbox{\tiny cur}}\leftarrow 2w_{\mbox{\tiny cur}}$
@@
 Algorithm 1 Aim-High
-3 Algorithm Specification
+3. Algorithm Specification
 The pseudocode for AH is given above, and each packet executes this algorithm upon being activated. Under AH, each packet begins with an initial window size of $w_{0}$ $2^{\mathcal{C}^{\epsilon}}$.
@@
 If a success occurs, the execution terminates; otherwise, the window doubles and another sample is executed.
-This process continues until $2^{j}$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
-3.1 Context and Overview
+This process continues until $j$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
+3.1. Context and Overview
 We start by considering how prior approaches fare in a model when there is a collision cost. Then, we discuss our design choices regarding AH, along with an overview of our analysis.
-3.1.1 A Baseline via Backoff.
+3.1.1. A Baseline via Backoff.
 A natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$.
 At $i=\lg(n)$, the probability of a success in a slot becomes constant.
-An example of its use in solving wakeup is given by Jurdziński and Stachowiak [14], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
+An example of its use in solving wakeup is given by Jurdziński and Stachowiak (Jurdziński and Stachowiak, 2002), who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
 So, we can observe that while backoff may be an important part of any solution, a key obstacle is how to set the initial sending probability suitably low.
@@
 What can we do?
-3.1.2 Design Choices and Overview.
+3.1.2. Design Choices and Overview.
 We now discuss how the design of AH allows us to do better than prior results in a model with collision cost.
@@
 At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots.
-This observation dictates the size of our samples, which are $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ and appear in lines 1 and 1 in the halving phase, and lines 1 and 1 in the doubling phase of our pseudocode.
+This observation dictates the size of our samples.
 In designing AH, we hope that $w^{*}\leq w_{0}$, since in this case, the above reasoning yields good expected collision cost (and, as we will show, good latency).
@@
 This choice sets up our analysis of the following two cases:
-  1. 1.
+  1. (1)
     If $n\sqrt{\mathcal{C}}\leq w_{0}$, then the good-window size is less than or equal to the initial-window size.
@@
 Of course, there is a tension with latency, since this adjustment cannot be too conservative, otherwise the latency will be too large.
-  2. 2.
+  2. (2)
     Else $n\sqrt{\mathcal{C}}>w_{0}$, and the good-window size is larger than the initial-window size.
@@
 Suppose we are in Case 1, but this failure occurs in the halving phase.
-Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 3).
+Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 4.3).
 However, if the halving phase is never revisited, the packets may accrue intolerable latency by executing the doubling phase without success (as the sending probability drops off exponentially fast).
@@
 Then, if we never reset $w_{\mbox{\tiny cur}}$ to (roughly) $w_{0}$, the latency costs may again be high.
-Therefore, iterations facilitate good expected latency (see Lemma 6).
+Therefore, iterations facilitate good expected latency (see Lemma 4.6).
 Parameter $\epsilon$.
@@
 However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$.
-Thus, degree of the polynomial increases with the reciprocal of $\epsilon$.
-4 Upper Bound
-In this section, we analyze the performance of AH.
+Thus, the degree of the polynomial increases with the reciprocal of $\epsilon$.
+4. Upper Bound
+In this section, we analyze the performance of AH; due to space constraints, we omit several proofs.
 Define $\eta$ $\max\{n,\mathcal{C}\}$.
 A good window was loosely discussed in Section 3.1.2; we define it as any window with size in $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$.
-Throughout, we assume that $n$ is sufficiently large.
-We make use of the following well-known facts.
-Fact 1
-The following inequalities hold.
-  1. (a)
-    For any $x$, $1-x\leq e^{-x}$,
-  2. (b)
-    For any $0\leq x\leq 1/2$, $1-x\geq e^{-2x}$.
-We also make use of the following standard Chernoff bounds.
-Theorem 4.1
-(Goemans [12]) Let $X=\sum_{i}X_{i}$, where $X_{i}$ are Bernoulli indicator random variables with probability $p$. The following holds:
-$Pr\left(X\geq(1+\delta)E[X]\right)$ $\leq\exp\left(-\frac{\delta^{2}E[X]}{2+\delta}\right)\mbox{~~~for any $\delta>0$}$
-and
-[…]
 The next lemma argues that success is likely in a good window.
-Lemma 1
+Lemma 4.1.
 W.h.p. in $\eta$, a success occurs if AH executes in a good window.
-Proof
-Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3). The probability of a success in any fixed slot of this good window is at least:
-[…]
-$\geq\left(\frac{1}{3e\sqrt{\mathcal{C}}}\right)\mbox{~~~since $\mathcal{C}\geq 4$.}$
-Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$.
-The expected number of successes in the sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots is at least:
-$E[X]\geq\left(\frac{d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})}{3e\sqrt{\mathcal{C}}}\right)$ $=\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$
-where $d^{\prime}$ is a constant that can be made arbitrarily large depending on $d$.
-We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
-$Pr\left(X\leq(1-\delta)\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$ $\leq Pr\left(X\leq\frac{d\ln(w_{\mbox{\tiny cur}})}{6e}\right)$
-$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
-$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
-$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
-$\leq\eta^{-d^{\prime\prime}}$
-where the third line follows from $w_{\mbox{\tiny cur}}=\Theta(\ln n\sqrt{\mathcal{C}})$, and $d^{\prime}$ and $d^{\prime\prime}$ can be arbitrarily large constants for $d$ is sufficiently large.
-4.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
+4.1. Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
 In this case, the good window size is less than or equal to $w_{0}$, so we argue that AH is likely to obtain a success during the halving phase.
 We make use of the following known result that upper bounds the probability of a collision.
-Lemma 2
+Lemma 4.2.
 (Biswas et al.
-[4]) Consider any slot $t$ where there exist $m$ active packets, each sending with the same probability $p$, where $p<1/m$.
+(Biswas et al., 2024a)) Consider any slot $t$ where there exist $m$ active packets, each sending with the same probability $p$, where $p<1/m$.
 The probability of a collision in slot $t$ is at most $\frac{2m^{2}p^{2}}{\left(1-mp\right)}$.
-Lemma 3
+Lemma 4.3.
 Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected total cost of collisions for AH is $O(\mathcal{C}^{(1/2)+\epsilon})$.
-Proof
+Proof.
 […]
 […]
-$G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ (1)
+(1) $G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$
 Upper Bounding $B_{j}$.
 Clearly, $B_{j}$ is upper-bounded by the cost of having a collision in every single slot of iterations $1$ through $j$.
-In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\ln(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contributes $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots.
-Additionally, in iteration $k$, the corresponding doubling phase executes $2^{k}$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, by the sum of a geometric series, over iterations $1$ through $j$, the doubling phases contributing $O(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j})$ slots.
+In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\lg(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contribute $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots.
+Additionally, in iteration $k$, the corresponding doubling phase executes $k$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, the doubling phases contribute $O(j^{2}\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}}))$ slots.
 By the above:
-$B_{j}$ $=O\left(j\,d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j}\right)\cdot\mathcal{C}$
-$=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\log(w_{0})2^{j}\right)\cdot\mathcal{C}$
-$=O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
-since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(2^{j}w_{0})=O(j\mathcal{C}^{\epsilon})$.
-By Lemma 1, the $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ where we can make the constant $c>0$ as large as we wish. Thus,
-$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
+$B_{j}$ $=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+j^{2}\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\cdot\mathcal{C}$
+$=O\left(j^{2}\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})\right)\cdot\mathcal{C}$
+$=O\left(j^{3}C^{(3/2)+2\epsilon}\right)$
+since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(jw_{0})=O(j\mathcal{C}^{\epsilon})$.
+By Lemma 4.1, $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ and we can make the constant $c>0$ as large as we wish. Thus,
+$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(j^{3}C^{(3/2)+2\epsilon})$
 $=O(1)$
@@
 Recall that the windows of this first halving phase are repeatedly halved.
-For a slot in a window of size $w=w_{0}/2^{k}$, for a positive integer $k\geq 0$, by Lemma 2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$), we have:
-$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}$
-$=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}$
+In a window of size $w=w_{0}/2^{k}$ where $k\geq 0$ is a positive integer, by Lemma 4.2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$):
+$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}.$
 The expected cost due to collisions that occur in this particular window is thus at most:
@@
 Note that the number of windows executed starting from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to size $n\sqrt{\mathcal{C}}$, is at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$; for ease of presentation, let $x=C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$.
-By the above, we can derive:
+We have:
 $G$ $\leq\sum_{k=0}^{x}\hskip 0.0pt\frac{2\left(\frac{n}{w_{0}/2^{k}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{k}}\right)\right)}\left(d\mathcal{C}^{(3/2)+\epsilon}\right)$
 $=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\sum_{k=0}^{x}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
-$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$ (2)
-$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$ (3)
-$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$ (4)
-$<\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{8}{\mathcal{C}}$
+(2) $\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$
+(3) $\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$
+(4) $=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$
 $=O\left(\mathcal{C}^{(1/2)+\epsilon}\right).$
@@
 Thus, Equation 1 is bounded as claimed.
-The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears.
-Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
-Lemma 4
+∎
+The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 4.3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears.
+Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
+Lemma 4.4.
 Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
-4.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
+4.2. Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
 In this case, the good window size exceeds $w_{0}$, and so we argue that AH likely succeeds during the doubling phase. A key insight is that, for Case 2, $\mathcal{C}$ is small in terms of $n$.
-Lemma 5
+Lemma 4.5.
 Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, $\mathcal{C}=O(\log^{1/\epsilon}n)$.
-Proof
+Proof.
 Since $n>2^{\mathcal{C}^{\epsilon}}/\sqrt{\mathcal{C}}$, we have:
 $\lg n$ $>\mathcal{C}^{\epsilon}-(\lg\mathcal{C})/2$
-$>\mathcal{C}^{\epsilon}-(\mathcal{C}^{\epsilon}/2).$ (5)
-To justify the second line, let $\delta\in(0,1)$, and note that:
-$(\lg\mathcal{C})/2$ $<\mathcal{C}^{\epsilon}/2$
-so long as $\epsilon>\frac{\lg\lg(\mathcal{C})}{\lg\mathcal{C}}$, which holds by assumption since $\epsilon$ is a constant, and this is well-defined since $\mathcal{C}\geq 4$.
-Returning to Equation 5, we have that:
-$(2\lg n)^{1/\epsilon}$ $>\mathcal{C}.$
-as desired, since $\epsilon$ is a constant.
+(5) $>\mathcal{C}^{\epsilon}-(\mathcal{C}^{\epsilon}/2).$
+where the second line holds so long as $\epsilon>\frac{\lg\lg(\mathcal{C})}{\lg\mathcal{C}}$, which is true by assumption since $\epsilon$ is a constant, and this is well-defined since $\mathcal{C}\geq 4$.
+Returning to Equation 5, we have that $(2\lg n)^{1/\epsilon}>\mathcal{C}$, as desired, since $\epsilon$ is a constant.
+∎
 The next lemma establishes bounds on expected latency and total collision cost, and the intuition is as follows.
-In Case 2, collisions are “cheap” by Lemma 5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$.
-Thus, the cost of collisions incurred until the good window is reached is small; it takes only $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=O(\log n)$ doublings of the window size to reach a good window in the doubling phase.
-Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by $j=O(\log\log n)$.
+In Case 2, collisions are “cheap” by Lemma 4.5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$.
+Thus, the cost of collisions incurred until the good window is reached is small; it takes only $O(\log n)$ doublings in the doubling phase.
 After this point the probability of success is high, and so subsequent iterations contribute little to the expected total cost of collisions.
-Lemma 6
-Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+4}n)\right)$.
-Proof
-The number of slots incurred for each individual halving phase is the sample size times the number of halvings, so upper bounded by:
-$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(2^{j}w_{0})\log(w_{0})\right)$
-$=O\left(\sqrt{\mathcal{C}}j\mathcal{C}^{2\epsilon}\right)$
-$=O\left(j\log^{1/(2\epsilon)+2}n\right)$
-where the last line follows by applying Lemma 5. Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
-The number of slots in the doubling phase of the $j$-th iteration is:
-$2^{j}d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O(2^{j}\sqrt{\mathcal{C}}\log(2^{j}w_{0}))$
-$=O(2^{j}\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon})$
-$=O(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n))$
-slots, where the last line follow by Lemma 5. By the above, summing up the number of slots executed over iterations 1 to $j$ is $O(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n))$.
+Lemma 4.6.
+Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+5}n)\right)$.
+Proof.
+The number of slots incurred for each individual halving phase is the sample size times the number of halvings, which is:
+$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(w_{0})\log(w_{0})\right)$
+$=O\left(\sqrt{\mathcal{C}}\mathcal{C}^{2\epsilon}\right)$
+$=O\left(\log^{1/(2\epsilon)+2}n\right)$
+where the second line follows from $w_{0}=2^{\mathcal{C}^{\epsilon}}$, and the last line follows by applying Lemma 4.5.
+Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
+The number of slots in the doubling phase of the $j$-th iteration is at most:
+$jd\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O\left(j\sqrt{\mathcal{C}}\log(2^{j}w_{0})\right)$
+$=O\left(j\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon}\right)$
+$=O\left(j^{2}\,(\log^{1/(2\epsilon)+1}n)\right)$
+slots, where the last line follow by Lemma 4.5. By the above, summing up the number of slots executed in the doubling phases over iterations 1 to $j$ is $O(j^{3}\,(\log^{1/(2\epsilon)+2}n))$.
 Pessimistically, let us assume no success ever occurs during the halving phase.
 Focusing on the doubling phase, a good window is reached within at most $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=\log_{2}n+O(\log\log n)$ doublings of the window size.
-Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log\log n$, where $c^{\prime}>0$ is a sufficiently large constant.
+Since $j$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log n$, where $c^{\prime}>0$ is a sufficiently large constant.
 By the above, the total number of slots up until a good window is reached is:
-$O\left(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log n)\,(\log\log n)^{2}(\log^{1/(2\epsilon)+2}n)\right)$
-$=O\left(\log^{1/(2\epsilon)+4}n\right)$
-After this point the probability of success is high, by Lemma 1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
-$O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
-$=O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{n}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-10.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
-$=O\left(\log^{1/(2\epsilon)+4}n\right)$
-which completes the argument.
-The following lemma about the expected total cost of collisions is proved similarly to Lemma 6.
-The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
-Lemma 7
-Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+4}n\right)$.
-4.3 Performance of AH
-We can now prove our upper bounds on the latency and expected collision cost for AH.
+$O\left(j^{3}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log^{3}n)(\log^{1/(2\epsilon)+2}n)\right)$
+$=O\left(\log^{1/(2\epsilon)+5}n\right)$
+After this point the probability of success is high, by Lemma 4.1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
+$O\left(\log^{1/(2\epsilon)+5}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(j^{3}\,(\log^{1/(2\epsilon)+2}n)\right)$
+$=O\left(\log^{1/(2\epsilon)+5}n\right)$
+which completes the argument. ∎
+The following lemma is proved similarly to Lemma 4.6. The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
+Lemma 4.7.
+Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+5}n\right)$.
+4.3. Performance of AH
+We can now prove our upper bound.
 Theorem 1.1. AH solves the wakeup problem with the following costs:
-  1. 1.
+  1. (1)
     If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
-  2. 2.
-    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
-Proof
-The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are there are two cases to consider.
+  2. (2)
+    Else, the expected latency is $O(\log^{5+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{5+\frac{3}{2\epsilon}}n)$.
+Proof.
+The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are two cases to consider.
 In Case 1, $n\sqrt{\mathcal{C}}\leq w_{0}$.
-By Lemmas 3 and 4, the expected latency and the expected total collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
+By Lemmas 4.3 and 4.4, the expected latency and the expected total collision cost for AH are respectively $O(\mathcal{C}^{(1/2)+2\epsilon})$ and $O(\mathcal{C}^{(1/2)+\epsilon})$, yielding (1).
 In the second case, $n\sqrt{\mathcal{C}}>w_{0}$.
-By Lemmas, 6 and 7, the expected latency and expected total cost of collisions are $O(\log^{4+\frac{1}{2\epsilon}}n)$ and $O(\log^{4+\frac{3}{2\epsilon})}n)$, respectively, yielding (2).
-5 Lower Bound
+By Lemmas, 4.6 and 4.7, the expected latency and expected total cost of collisions are $O(\log^{5+\frac{1}{2\epsilon}}n)$ and $O(\log^{5+\frac{3}{2\epsilon}}n)$, respectively, yielding (2).
+∎
+5. Lower Bound
 […]
-Lemma 8
+Lemma 5.1.
 Any fair algorithm for the wakeup problem that succeeds w.h.p. in $\eta$ must have total contention at least $2$.
-Proof
 […]
-$TC(\mathcal{A})=\sum_{u}LC(u)=n\cdot LC(u)$
-By our assumption, $n\cdot LC(u)<2$.
-The probability of $u$ not succeeding in slot any fixed $t$ is at least 1p(t)1-p(t).
-Therefore, the probability that $u$ never succeeds over the execution of $\mathcal{A}$ is:
-$\geq\prod_{i=1}^{\lambda}(1-p(i))$
-$\geq e^{-2\sum_{i}^{\lambda}p(i)}$
-$\geq e^{-2\cdot LC(u)}$
-$\geq e^{-2\cdot TC(\mathcal{A})/n}$
-$>e^{-4/n}$
-where the second line follows from Fact 1(b), and the third line follows from the definition of $LC(u)$. The last line follows from the assumption that $TC(\mathcal{A})<2$.
-It follows that the probability of all packets failing over the execution of $\mathcal{A}$ is:
-$\geq\left(e^{-4/n}\right)^{n}=\frac{1}{e^{4}}$
-Thus, there is a constant probability of all packets failing, which contradicts the w.h.p. guarantee of $\mathcal{A}$.
-[…]
-Lemma 9(Biswas et al. [5], Lemma 24)
+Lemma 5.2 (Biswas et al. (Biswas et al., 2024b), Lemma 24).
 Fix any slot $t$ in any fair algorithm for the wakeup problem. If $\texttt{Con}(t)>2$, then the probability of a collision is at least $1/10$.
-By Lemma 9, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$.
+By Lemma 5.2, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$.
 Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$.
@@
 We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
-Lemma 10
+Lemma 5.3.
 Fix any slot $t$ where there are at least $2$ packets and $\texttt{Con}(t)\leq 2$ under the execution of any fair algorithm for the wakeup problem.
 The probability of a collision in $t$ is $\Omega(\texttt{Con}^{2}(t))$.
-Proof
-For a fair algorithm, in any slot $t$, $\texttt{Con}(t)=np$.
-For any fixed slot, the probability of a collision involving exactly two packets (since there are at least $2$ packets by assumption) is a lower bound on the probability of a collision (that may involve more than $2$ packets), and this is $\binom{n}{2}p^{2}(1-p)^{n-2}$, which is at least:
-$\geq\binom{n}{2}p^{2}(1-p)^{n-2}$
-$\geq\binom{n}{2}p^{2}(1-p)^{n}\geq\left(\frac{n^{2}}{2}\right)\left(\frac{\texttt{Con}(t)}{n}\right)^{2}e^{-2\cdot\texttt{Con}(t)}\geq\left(\frac{\texttt{Con}^{2}(t)}{2e^{2\cdot\texttt{Con}(t)}}\right)$
-$=\Omega(\texttt{Con}^{2}(t))$
-where the last line follows from $\texttt{Con}(t)\leq 2$.
 We can derive a relationship between latency and expected collision cost for fair algorithms.
@@
 Any fair algorithm for the wakeup problem that has $O(\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n)$ latency has an expected total cost of collisions $\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$.
-Proof
+Proof.
 Let $\mathcal{A}$ be an algorithm that w.h.p.
-in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is any constant.
-By Lemma 8, $\mathcal{A}$ has total contention:
-$\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$ (6)
-As we discussed, an implication of Lemma 9, we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot.
-Thus, we can use Lemma 10 (noting that $n\geq 2$), which implies that the expected cost of collisions in a single slot $t$ is:
+in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is some constant.
+By Lemma 5.1, $\mathcal{A}$ has total contention:
+(6) $\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$
+As we discussed, Lemma 5.2 implies that we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot.
+Thus, we can use Lemma 5.3 (noting that $n\geq 2$), which implies that the expected cost of collisions in a single slot $t$ is:
 […]
 for some constant $k>0$. So, the expected total cost of collisions is:
-$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\mathcal{C}$
-By Jensen’s inequality for convex functions, we have:
-$\frac{\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)}{L}\geq\left(\frac{\sum_{t=1}^{L}k\,\texttt{Con}(t)}{L}\right)^{2}$
-which implies:
-$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\geq\frac{\left(\sum_{t=1}^{L}k\,\texttt{Con}(t)\right)^{2}}{L}$ (7)
-So, by Equation 7, the expected cost is at least:
-[…]
-and since $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$, we have:
-$\frac{\mathcal{C}\,k^{2}\,2^{2}}{L}$ $=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
-which yields the claim.
+$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\,\mathcal{C}.$
+Using Jensen’s inequality for convex functions, we have:
+$\mathcal{C}k\sum_{t=1}^{L}\texttt{Con}^{2}(t)$ $\geq\mathcal{C}k\frac{\left(\sum_{t=1}^{L}\texttt{Con}(t)\right)^{2}}{L}$
+$\geq\frac{\mathcal{C}\,k\,2^{2}}{L}$
+$=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
+where the second line follows from Equation 6, and the last line follows since $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$. ∎
 As evidence that our result is not far off from the asymptotic optimal, we note that any fair algorithms with latency $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ must have an expected total cost of collisions that is $\Omega(\mathcal{C}^{\frac{1}{2}-2\epsilon})$, which is an $O(\mathcal{C}^{3\epsilon})$-factor smaller than that achieved by AH.
-6 Conclusion and Future Work
-The theoretical exploration of collision costs remains in its early stages. Our work here provides steps towards addressing fundamental problems that connect with the wireless domain.
-For future work, we aim to prove more general lower bounds. Additionally, we wish to derive results for the dynamic setting, where packets arrive over time rather than all being initially present.
+6. Conclusion and Future Work
+Our work gives a preliminary theoretical exploration addressing collision costs.
+For future work, we aim to prove more general lower bounds.
+We also wish to derive results for the dynamic setting, where packets arrive over time rather than all being initially present.
+Acknowledgements. This research is supported by the National Science Foundation grants CNS-2210300 and CCF-2144410.